### v1

# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

Abstract
Signed networks, where edges are labeled as positive or negative to indicate friendly or antagonistic interactions, offer a natural framework for studying polarization, trust, and conflict in social systems. Detecting meaningful group structures in these networks is crucial for understanding online discourse, political division, and trust dynamics. A key challenge is to identify groups that are cohesive internally yet antagonistic externally, while allowing for neutral or unaligned vertices. In this paper, we address this problem by identifying $k$ polarized communities that are large, dense, and balanced in size. We develop an approach based on Frank-Wolfe optimization, leading to a local search procedure with provable convergence guarantees. Our method is both scalable and efficient, outperforming state-of-the-art baselines in solution quality while remaining competitive in terms of computational efficiency.
1 Introduction
Signed networks extend traditional graph representations by associating each edge with a positive or negative number, indicating friendly or antagonistic relationships. Originating from studies on social dynamics in the 1950s (Harary, 1953), signed networks introduce fundamental differences in graph structure that make many algorithms designed for unsigned networks inapplicable (Tang et al., 2016; Bonchi et al., 2019; Tzeng et al., 2020). These challenges have fueled extensive research in recent years, leading to advances in signed network embeddings, signed clustering, and signed link prediction. We refer to the survey by (Tang et al., 2016) for a comprehensive review of these methods. Most relevant to this paper is the problem of signed clustering, which we split into two categories: (i) signed network partitioning (SNP), and (ii) polarized community discovery (PCD). The latter is the problem studied in this paper.
The goal of signed clustering is to identify $k$ conflicting groups (clusters) where intra-cluster similarity is maximized (predominantly positive) and inter-cluster similarity is minimized (predominantly negative). This problem has numerous real-world applications (Tang et al., 2016), particularly in social networks, where vertices represent individuals and edges capture friendly or antagonistic relationships (e.g., shared or opposing political views). Detecting conflicting groups in such networks is crucial for analyzing polarization (Adamic and Glance, 2005; Yardi and Boyd, 2010; Xiao et al., 2020), echo chambers (Garrett, 2009; Flaxman et al., 2016), and the spread of misinformation (Shu et al., 2017; Cooke, 2018; Yang et al., 2019).
In the SNP problem, the $k$ groups must form a partition of the vertices, meaning every vertex must be included. Spectral methods based on the signed Laplacian have been widely used to tackle this problem (Kunegis et al., 2010; Chiang et al., 2012; Mercado et al., 2019; Cucuringu et al., 2019). Alternatively, formulating SNP explicitly as an optimization problem leads to the well-studied correlation clustering (CC) problem (Bansal et al., 2004), which is known to be APX-hard. Consequently, numerous approximation algorithms have been developed (Bansal et al., 2004; Charikar et al., 2005; Demaine et al., 2006; Ailon et al., 2008), with local search methods standing out for their strong performance in both clustering quality and computational efficiency (Thiel et al., 2019; Chehreghani, 2023; Aronsson and Chehreghani, 2024a; b).
The problem formulation of PCD is identical to that of SNP, except that the $k$ clusters are not required to form a partition of the vertices, allowing some vertices to remain unassigned. The goal is therefore to only find the dense subgraphs of polarized communities. This accounts for cases where certain vertices are neutral w.r.t. the underlying conflicting group structure. For example, in a social network with a heated political debate, many users may not engage in the dispute, and their interactions might not align with any specific faction. There is a substantial body of work addressing this problem, but most approaches focus on identifying only two communities (Bonchi et al., 2019; Xiao et al., 2020; Ordozgoiti et al., 2020; Fazzone et al., 2022; Niu and Sariyüce, 2023; Gullo et al., 2024). As a result, they do not easily generalize to arbitrary $k$. To our knowledge, only two works specifically tackle PCD for arbitrary $k$. (Chu et al., 2016) formulated the task as a constrained quadratic optimization problem and proposes an efficient algorithm that iteratively refines small subgraphs, avoiding the costly computation of the full adjacency matrix. (Tzeng et al., 2020) introduced a spectral method based on maximizing a discrete Rayleigh quotient, which extends the seminal work of (Bonchi et al., 2019) to accommodate arbitrary $k$. Their method is known to produce highly imbalanced communities in terms of size (Gullo et al., 2024).
The main contributions of this paper are as follows: (i) We propose a novel formulation for the PCD problem that encourages more balanced communities, addressing a key limitation of (Tzeng et al., 2020). (ii) We demonstrate that this formulation is particularly well-suited for optimization via the block-coordinate Frank-Wolfe algorithm (Frank and Wolfe, 1956; Lacoste-Julien et al., 2013). (iii) We establish a simple yet effective equivalent local search method with a provable convergence rate. (iv) We propose techniques that allow the local search method to scale to large networks. (v) Finally, through extensive experiments, we show that our approach consistently outperforms state-of-the-art baselines.
2 Problem Formulation
We start by introducing the relevant notation, followed by an introduction to correlation clustering (CC), which is connected to our problem. Finally, we present the problem of polarized community discovery (PCD).
2.1 Notation
Consider a signed network $G=(V,E)$, where $V$ is the set of objects and $E$ the set of edges. The weight of an edge $(i,j)\in E$ is represented by the element $A_{i,j}\in\{-1,0,+1\}$ of an adjacency matrix $A$. The matrix $A$ is symmetric with zeros on the diagonal, which means $A_{i,j}=A_{j,i}$ and $A_{i,i}=0$. We use $A_{i,:}$ and $A_{:,j}$ to denote row $i$ and column $j$ of $A$, respectively. While we restrict all similarities to be in $\{-1,0,+1\}$ (for clarity), all methods presented in the paper extend to arbitrary similarities in $\mathbb{R}$. We can decompose the adjacency matrix as $A=A^{+}-A^{-}$ where $A^{+}=\max(A,0)$ and $A^{-}=\max(-A,0)$. A clustering with $k$ clusters is denoted $S_{[k]}=\{S_{1},\dots,S_{k}\}$, where each $S_{m}\subseteq V$ is the set of objects assigned to cluster $m\in[k]=\{1,\dots,k\}$. Let $N^{+}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{+}_{i,j}$ and $N^{-}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{-}_{i,j}$ be the sum of positive and absolute negative intra-cluster similarities, respectively. Furthermore, let $N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A^{+}_{i,j}$ and $N^{-}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A^{-}_{i,j}$ be the sum of positive and absolute negative inter-cluster similarities, respectively.
2.2 Correlation Clustering
We begin by noting that for CC, unlike PCD to be discussed in the next subsection, a clustering $S_{[k]}$ is a partition of $V$, meaning $V=\bigcup_{m\in[k]}S_{m}$ and each $S_{m}$ is disjoint. A notable feature of CC is its ability to automatically determine the number of clusters based on the similarities encoded in $A$. However, a variant of this problem has also been studied where the number of clusters, $k$, is specified as an input (Giotis and Guruswami, 2006). Automatic detection of the number of clusters could be a desirable property of a clustering algorithm. However, constraining the number of clusters to $k$ can act as a form of regularization, and has been shown to produce higher-quality clusters in many scenarios (Chehreghani et al., 2012). Given this, the $k$-CC problem can be defined as shown below.
Problem 1 ($k$-CC).
Find a clustering $S_{[k]}$ that maximizes
$N^{+}_{\text{intra}}-N^{-}_{\text{intra}}+N^{-}_{\text{inter}}-N^{+}_{\text{inter}}.$ (1)
In other words, we want to find a clustering that maximizes the sum of intra-cluster similarities and minimizes the sum of inter-cluster similarities. The following proposition presents alternative objectives equivalent to maximizing Eq. 1. While this is known in the CC literature (Chehreghani, 2013), we include a complete summary here to better contextualize our problem.
Proposition 1.
Problem 1 is equivalent to finding a clustering $S_{[k]}$ that maximizes any one of the four objectives below [^1] By equivalent, we mean they share all local maxima, including the global maximum..
$N^{+}_{\text{intra}}+N^{-}_{\text{inter}}$ (2)
$-N^{-}_{\text{intra}}-N^{+}_{\text{inter}}$ (3)
$N^{+}_{\text{intra}}-N^{-}_{\text{intra}}$ (4)
$N^{-}_{\text{inter}}-N^{+}_{\text{inter}}$ (5)
Furthermore, maximizing any other combination of the four terms is not equivalent to Problem 1.
All proofs can be found in Appendix A. CC is typically formulated as either maximizing agreements (Eq. 2) or minimizing disagreements (Eq. 3) (Bonchi et al., 2014). It can also be framed in terms of maximizing intra-cluster similarities, referred to as max correlation (Eq. 4), or minimizing inter-cluster similarities (Eq. 5). Finally, we can combine all these notions into one single objective (i.e., Eq. 1). CC is an NP-hard problem, leading to the development of numerous approximation algorithms. Existing approximation algorithms maximize one of the five expressions above, leading to differences in clustering performance, computational complexity and theoretical performance guarantees.
2.3 Polarized Community Discovery
The problem of PCD is similar to CC in that the goal is to identify $k$ clusters $S_{1},\dots,S_{k}$, where the similarity within each cluster is large (and positive) and the similarity between different clusters is small (and negative). However, unlike CC, we also allow objects to remain neutral, introducing an additional neutral set $S_{0}$. This means an object can either be assigned to one of the non-neutral clusters $S_{1},\dots,S_{k}$ or designated as neutral by assigning it to $S_{0}$. Assigning an object to $S_{0}$ effectively excludes it from influencing the objective function. Specifically, assigning an object $i\in V$ to $S_{0}$ is equivalent to setting both row $A_{i,:}$ and column $A_{:,i}$ of the matrix $A$ to zero. Consequently, a clustering $S_{[k]}$ is no longer a partition of $V$ and we have $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ (although all clusters are still disjoint). In our view, the goal of PCD is to extract $k$ non-neutral clusters that are (i) large, (ii) balanced, and (iii) dense. Dense clusters imply that each object should strongly align with its assigned cluster—being highly similar to most objects within its cluster and markedly dissimilar to those in other clusters. Objects lacking a clear association, such as those similar to objects in multiple clusters or those that are inherently neutral (e.g., low-degree objects), should be labeled as neutral by assigning them to $S_{0}$. We notice that a natural trade-off exists between the size of the non-neutral clusters and their density, as very small clusters can achieve high density trivially.
We begin this section by explaining why Eq. 1, which incorporates all relevant terms, must be considered when neutral objects are allowed. Much prior work on PCD also optimize all terms, but often without providing a detailed justification for this choice. The next proposition provides such an intuition.
Proposition 2.
A clustering $S_{[k]}$ with neutral objects $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ that maximizes one of the objectives in Eqs. 1-5 is not guaranteed to maximize any of the other objectives [^2] Unless $k=2$, in which case Eq. 2 and Eq. 1 are equivalent as established in (Bonchi et al., 2019)..
From Proposition 2, we conclude that each term in Eq. 1 provides unique information when neutral objects are allowed, unlike the standard CC problem, where the different objectives are equivalent, as outlined by Proposition 1. This makes Eq. 1 the most reasonable objective for optimization in this context, as it effectively balances all contributing terms. Moreover, since each term captures unique aspects of the PCD problem, it may be beneficial to weight them differently to achieve an optimal trade-off. Furthermore, (Bonchi et al., 2019) showed that for $k=2$, an optimal solution to Eq. 1 consists of no neutral objects, i.e., $S_{0}=\emptyset$. While we cannot directly extend this conclusion to $k>2$ (see proof of Proposition 2), it is evident that even for $k>2$, an object may be assigned to a non-neutral cluster as long as it marginally improves the objective in Eq. 1, even if it does not maintain the density of the graph induced by non-neutral clusters $S_{[k]}$. In other words, low-degree objects that should ideally remain neutral may be included in non-neutral clusters.
Based on the above discussion, we conclude that (i) it may be beneficial to weight the terms in Eq. 1 differently and (ii) we must encourage the presence of neutral objects by penalizing large/sparse non-neutral clusters. In prior work, these concerns are typically addressed by weighting inter-cluster terms with a parameter $\alpha\in\mathbb{R}$ and normalizing Eq. 1 by the number of non-neutral objects, i.e.,
$\frac{(N^{+}_{\text{intra}}-N^{-}_{\text{intra}})+\alpha(N^{-}_{\text{inter}}-N^{+}_{\text{inter}})}{\sum_{m\in[k]}|S_{m}|}.$ (6)
If $\alpha=1/(k-1)$, Eq. 6 is commonly referred to as polarity in prior work and is a well-established objective for PCD (Bonchi et al., 2019; Tzeng et al., 2020). However, as highlighted in (Gullo et al., 2024), maximizing polarity often results in highly imbalanced clustering solutions. In particular, clustering solutions with the same polarity can differ significantly in terms of cluster balance. A concrete example illustrating this issue is provided in Appendix C.
In this paper, we propose an alternative objective that, instead of normalizing by the number of non-neutral objects, incorporates a regularization term by subtracting the sum of squared sizes of the non-neutral clusters [^3] For simplicity, we follow prior work by only weighting the inter-cluster terms.:
$(N^{+}_{\text{intra}}-N^{-}_{\text{intra}})+\alpha(N^{-}_{\text{inter}}-N^{+}_{\text{inter}})-\beta\sum_{m\in[k]}|S_{m}|^{2}.$ (7)
The third term in Eq. 7 has been previously applied to the minimum cut objective for unsigned networks (Chehreghani, 2023) where in practice, however, they use a different penalty: they shift the pairwise similarities so that the sum of the rows and columns of the similarity matrix becomes zero. In our context (i.e., for the PCD problem), this objective achieves two goals simultaneously: it penalizes the formation of (i) large/sparse and (ii) imbalanced non-neutral clusters. The second property is easy to see, as for a clustering with $k$ clusters and $n$ objects, the term $\sum_{m\in[k]}|S_{m}|^{2}$ is minimized when each cluster is assigned $\frac{n}{k}$ objects (i.e., the clusters are perfectly balanced). The parameter $\beta\in\mathbb{R}$ allows us to easily trade-off the size of non-neutral clusters and their density, which is a desirable property in this context as discussed in the beginning of this section. As we will demonstrate in the subsequent sections, maximizing our objective with varying values of $\beta$ allows us to generate a range of clustering solutions with different cluster sizes and corresponding densities. This enables the selection of a solution that offers the most desirable trade-off, and is absent in the existing methods that are based on Eq. 6. Finally, we introduce regularization as an additive term rather than a normalization for three key reasons: (i) It allows for a flexible trade-off between cluster size and density. (ii) It enables the development of efficient optimization procedures with strong convergence guarantees. (iii) As we will demonstrate, it results in superior performance in practice. We are now ready to formally state our problem.
Problem 2 ($k$-PCD).
Find a clustering $S_{[k]}$ with neutral objects $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ that maximizes Eq. 7.
We start by highlighting the computational complexity of the problem with the following theorem.
Theorem 1.
Problem 2 (i.e., $k$-PCD) is NP-hard.
Theorem 1 underscores the necessity of approximate methods to solve Problem 2.
3 Algorithms
In this section, we demonstrate how our problem can be solved using Frank-Wolfe (FW) optimization (Frank and Wolfe, 1956). Specifically, we consider a variant called block-coordinate FW, which we begin by describing in the next subsection. After this, we establish its equivalence to a straightforward and provably efficient local search procedure. Next, we analyze the convergence rate of this approach. Following that, we propose practical enhancements to improve scalability, enabling the method to handle large problems. Finally, we provide a detailed analysis of the impact of $\alpha$ and $\beta$.
3.1 Block-Coordinate Frank-Wolfe Optimization
Algorithm 1 Block-coordinate Frank-Wolfe Optimization
1:  Initialize ${\bm{x}}_{[n]}^{(0)}\in\mathcal{D}^{(1)}\times\mathcal{D}^{(2)}\times\dots\times\mathcal{D}^{(n)}$.
2:  for $t\coloneqq 0,\dots,T$ do
3:     Select a random block $i\in[n]$
4:     ${\bm{x}}_{i}^{\ast}\coloneqq\operatorname*{arg\,max}_{{\bm{x}}_{i}\in\mathcal{D}^{(i)}}{\bm{x}}_{i}\cdot\nabla_{i}f({\bm{x}}_{[n]}^{(t)})$
5:     Let $\gamma\coloneqq\frac{2n}{t+2n}$ or optimize $\gamma$ by line-search
6:     ${\bm{x}}^{(t+1)}_{i}\coloneqq(1-\gamma){\bm{x}}^{(t)}_{i}+\gamma{\bm{x}}_{i}^{\ast}$
7:  end for
The Frank–Wolfe (FW) algorithm is one of the earliest methods for nonlinear constrained optimization (Frank and Wolfe, 1956). In recent years, it has regained popularity, particularly in machine learning, due to its scalability (Jaggi, 2013). In this paper, we use a variant of this method called block-coordinate FW (Lacoste-Julien et al., 2013). This method yields a significantly faster optimization procedure while enjoying similar theoretical guarantees. Block-coordinate FW is applied to problems where the feasible domain can be split into blocks $\mathcal{D}=\mathcal{D}^{(1)}\times\dots\times\mathcal{D}^{(n)}\subseteq\mathbb{R}^{d}$, where each $\mathcal{D}^{(i)}\subseteq\mathbb{R}^{d_{i}}$ is convex and compact and we have $d=\sum_{i=1}^{n}d_{i}$. Let ${\bm{x}}_{[n]}$ denote the concatenation of the variables ${\bm{x}}_{i}\in\mathcal{D}^{(i)}$ from all blocks $i\in[n]$. The optimization problem is then
$\max_{{\bm{x}}_{[n]}\in\mathcal{D}^{(1)}\times\dots\times\mathcal{D}^{(n)}}f({\bm{x}}_{[n]}),$ (8)
where $f$ is a differentiable function with an $L$-Lipschitz continuous gradient. This approach is particularly effective when optimizing $f$ w.r.t. the variables in a single block (while keeping other blocks fixed) is simple and efficient. This turns out to be the case for our problem, as will be discussed in the remainder of this section. The method is outlined in Alg. 1, where $\nabla_{i}f({\bm{x}}_{[n]})$ represents the gradient of $f({\bm{x}}_{[n]})$ with respect to block ${\bm{x}}_{i}$. When the problem involves only a single block ($n=1$), Alg. 1 reduces to the standard FW algorithm. We now show how our problem can be turned into an instance of Eq. 8. In the following proposition, we show an alternative way of writing our objective in Eq. 7.
Proposition 3.
Our objective in Eq. 7 can be written as
$\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}.$ (9)
We observe that the regularization term in Eq. 7 is equivalent to shifting the intra-cluster similarities by $-\beta$. This reformulation proves highly useful for the remainder of this section. We now describe how our problem is well-suited for block-coordinate FW. In our context, each object $i\in[n]$ defines a block. We represent the cluster membership of object $i$ using ${\bm{x}}_{i}\in\{{\bm{e}}_{0},\dots,{\bm{e}}_{k}\}$, where ${\bm{e}}_{m}$ (for $m\in\{0,\dots,k\}$) are the standard basis vectors. Each ${\bm{x}}_{i}$ is a vector of dimension $k+1$, with index zero indicating membership in the neutral set $S_{0}$. Specifically, if $x_{i0}=1$, object $i$ is assigned to $S_{0}$. Using this notation, we can now define our objective as follows.
$f({\bm{x}}_{[n]})=\sum_{(i,j)\in E}\sum_{m\in[k]}x_{im}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{(i,j)\in E}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}x_{im}x_{jp}A_{i,j}.$ (10)
Note that we do not include any terms involving $x_{i0}$, thereby excluding contributions from neutral objects, as intended. The objective in Eq. 10 remains discrete and is therefore unsuitable for FW optimization. To address this, we relax the problem to make it continuous by allowing soft cluster memberships. Specifically, each ${\bm{x}}_{i}\in\Delta^{k+1}$, where $\Delta^{k+1}=\{{\bm{x}}\in\mathbb{R}^{k+1}\mid x_{m}\geq 0,\sum_{m=0}^{k}x_{m}=1\}$ represents the simplex of dimension $k$. With this relaxation, we can now reformulate the optimization problem as follows.
$\max_{{\bm{x}}_{i}\in\Delta^{k+1},\forall i\in[n]}f({\bm{x}}_{[n]})$ (11)
Equation 11 is a specific instance of the block-coordinate FW formulation described in Eq. 8 (where $f$ is non-concave). Consequently, we can apply Alg. 1 to solve this problem.
3.2 Equivalence to a Local Search Approach
Algorithm 2 Local Search for PCD
1:  Randomly assign each object $i\in[n]$ to one of the clusters in $S_{[k]}$, or make it neutral by adding it to $S_{0}$
2:  while not converged do
3:     Select object $i\in[n]$ uniformly at random
4:     Assign object $i$ to a cluster (or make it neutral) which maximally increases our objective in Eq. 7.
5:  end while
We now show that optimizing Eq. 11 using Alg. 1 is equivalent to the local search procedure in Alg. 2. Let matrix $G\in\mathbb{R}^{n\times(k+1)}$, where element $G_{i,m}\triangleq[\nabla_{i}f({\bm{x}}^{(t)}_{[n]})]_{m}$ is the gradient of $f({\bm{x}}_{[n]})$ w.r.t. variable $m$ of block $i$ evaluated at ${\bm{x}}^{(t)}_{[n]}$ (solution at step $t$ of Alg. 1). Given this, we present the following theorem.
Theorem 2.
If ${\bm{x}}_{[n]}^{(0)}$ in Alg 1 is discrete, the following holds. (a) For our problem (Eq. 11), the solution ${\bm{x}}^{\ast}_{i}$ (line 4 of Alg. 1) is the basis vector ${\bm{e}}_{p}$, where $p=\operatorname*{arg\,max}_{m\in\{0,\dots,k\}}G_{i,m}$ and the optimal value of the step size on line 6 is $\gamma=1$. (b) Our objective function in Eq. 10 satisfies $({\bm{x}}^{\ast}_{i}-{\bm{x}}^{(t)}_{i})\cdot G_{i,:}=f({\bm{x}}^{\ast}_{[n]})-f({\bm{x}}_{[n]}^{(t)})$, where ${\bm{x}}^{\ast}_{[n]}$ is ${\bm{x}}^{(t)}_{[n]}$ with block $i$ modified to ${\bm{x}}^{\ast}_{i}$.
From part (a) of Theorem 2, the current solution, ${\bm{x}}^{(t)}_{[n]}$, remains discrete (i.e., hard cluster assignments) at every step of Alg. 1 for all $i\in[n]$. Moreover, each step of Alg. 1 consists of placing object $i$ in the cluster $m\in\{0,\dots,k\}$ with maximal gradient $G_{i,m}$. By part (b) of Theorem 2, this is equivalent to placing object $i$ in the cluster that maximally improves our objective in Eq. 7. Based on this, we conclude the following corollary.
Corollary 1.
From Theorem 2, if ${\bm{x}}_{[n]}^{(0)}$ is discrete, solving the optimization problem in Eq. 11 using Alg. 1 is equivalent to executing the local search procedure described in Alg. 2.
Given this, we now present results for the convergence rate of Alg. 2.
3.3 Convergence Analysis
Following the prior work on the analysis of general FW algorithms (Jaggi, 2013; Lacoste-Julien et al., 2013), we begin by providing the following definitions.
Definition 1 (FW duality gap).
The FW duality gap is defined as (Jaggi, 2013)
$g({\bm{x}}_{[n]})=\max_{{\bm{s}}_{[n]}\in\mathcal{D}}({\bm{s}}_{[n]}-{\bm{x}}_{[n]})\cdot\nabla f({\bm{x}}_{[n]}),$ (12)
which is zero if and only if ${\bm{x}}_{[n]}$ is a stationary point. Furthermore, let $\tilde{g}_{t}=\min_{0\leq l\leq t-1}g({\bm{x}}^{(l)}_{[n]})$ be the smallest duality gap observed in Alg. 1 up until step $t$.
Definition 2 (Convergence rate).
We say the convergence rate of Alg. 1 is at least $O(1/r_{t})$ if $\mathbb{E}[\tilde{g}_{t}]\leq O(1/r_{t})$, where $r_{t}$ is some expression involving only $t$ and the expectation is w.r.t. the random selection of blocks on line 3. If $n=1$ the bound is deterministic.
The FW algorithm has been shown to converge to a stationary point of $f$ under various settings, with well-established convergence rates. We summarize a few known results below. The standard FW algorithm ($n=1$) achieves a deterministic convergence rate of $O(1/t)$ for concave $f$ (Frank and Wolfe, 1956) and $O(1/\sqrt{t})$ for non-concave $f$ (Lacoste-Julien, 2016; Reddi et al., 2016). For the block variant, (Lacoste-Julien et al., 2013) prove a convergence rate of $O(1/t)$ for concave $f$ in expectation. For non-concave $f$, (Thiel et al., 2019) prove a convergence rate of $O(1/t)$ in expectation, under the assumption that $f({\bm{x}}_{[n]})$ is multilinear in each block ${\bm{x}}_{i}$ including correlation clustering. We here extend the analysis of (Thiel et al., 2019) to Problem 2 ($k$-PCD) using Alg. 2, described in Theorem 3. Note that their analysis cannot be applied directly to our objective function in Eq. 10 as this objective does not satisfy the multilinearity property.
Theorem 3.
The convergence rate of Alg. 2 is at least $nh_{0}/t=O(1/t)$, where $h_{0}=\sum_{(i,j)\in E}|A_{i,j}|$.
The $O(1/t)$ convergence rate presented in Theorem 3 should be compared with the deterministic convergence rate of $O(1/\sqrt{t})$ for general non-concave functions $f$ under the standard FW method ($n=1$) (Lacoste-Julien, 2016; Reddi et al., 2016).
3.4 Improving the Computational Complexity
Algorithm 3 Local Search for PCD (efficient)
1:  Randomly assign each object $i\in[n]$ to one of the clusters in $S_{[k]}$, or make it neutral by adding it to $S_{0}$
2:  Initialize $X$ based on initial clustering $S_{[k]}$
3:  $M\coloneqq 2AX$
4:  while not converged do
5:     Select object $i\in[n]$ uniformly at random
6:     $\hat{p}\coloneqq\text{current cluster of }i$
7:     $\eta_{i}\coloneqq\sum_{p}M_{i,p}$
8:     $G_{i,p}\coloneqq-\beta+(1+\alpha)M_{i,p}-2\beta|S_{p}|+2\beta\bm{1}_{[i\in S_{p}]}-\alpha\eta_{i},\forall p\in[k]$
9:     $G_{i,0}\coloneqq 0$
10:     $p^{\ast}\coloneqq\operatorname*{arg\,max}_{p\in\{0,\dots,k\}}G_{i,p}$
11:     If $p^{\ast}=\hat{p}$, then skip to next iteration
12:     Assign object $i$ to cluster $S_{p^{\ast}}$
13:     if $\hat{p}\in[k]$ (current cluster non-neutral) then
14:        $M_{:,\hat{p}}\coloneqq M_{:,\hat{p}}-2A_{:,i}$
15:     end if
16:     if $p^{\ast}\neq 0$ (new cluster not neutral) then
17:        $M_{:,p^{\ast}}\coloneqq M_{:,p^{\ast}}+2A_{:,i}$
18:     end if
19:  end while
In the previous section, we demonstrated that Alg. 2 is guaranteed to converge at the linear rate $O(1/t)$, making it highly efficient. In this section, we propose an alternative version of Alg. 2, designed to enhance the efficiency of each step $t$ while maintaining full equivalence in functionality. This ensures that the convergence analysis from the previous section still remains valid. Firstly, a naive implementation of Alg. 2 has a complexity of $O(Tk^{2}n^{2})$, as each iteration requires $O(k^{2}n^{2})$ to compute the full objective in Eq. 9 for every candidate cluster in order to determine the best cluster for the current object $i$. Since the number of iterations $T$ until convergence is typically larger than $n$, this approach can become computationally expensive.
Part (b) of Theorem 2 offers an alternative: instead of evaluating the full objective, we can compute the gradient $G_{i,:}$, which involves only terms related to object $i$. Let $M_{i,m}\triangleq 2\sum_{j\in S_{m}}A_{i,j}$, $\eta_{i}\triangleq\sum_{m\in[k]}M_{i,m}$ and $\beta_{im}=2\beta|S_{m}|-2\beta\bm{1}_{[i\in S_{m}]}$. Given this, we present the following theorem.
Theorem 4.
Let $S_{[k]}$ be the current clustering of our local search procedure, with neutral objects $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$. The gradient can then be expressed as follows.
$G_{i,m}=-\beta+(1+\alpha)M_{i,m}-\beta_{im}-\alpha\eta_{i}$ (13)
for all $m\in[k]$ and $G_{i,0}=0$.
A naive calculation of the full gradient $G_{i,:}$ for block $i$ is $O(k^{2}n)$. However, the particular form of the gradient presented in Eq. 13 makes it $O(kn)$. See the proof of Theorem 4 for further insight on this. From Theorem 2, the gradient $G_{i,m}$ represents the impact on the full objective in Eq. 9 if object $i$ is placed in cluster $m$. Thus, because $G_{i,0}=0$, we observe that an object $i$ is made neutral if its contribution to all non-neutral clusters is currently negative. Moreover, the total complexity is now reduced to $O(Tkn)$, which is a significant improvement over the naive approach with complexity of $O(Tk^{2}n^{2})$.
We present a third approach, outlined in Alg. 3, where we define a matrix $X\in\{0,1\}^{n\times k}$, with $X_{i,m}=1$ if object $i$ belongs to cluster $m\in[k]$, and zero otherwise. Neutral objects $i\in S_{0}$ have rows $X_{i,:}$ of zeros. The procedure precomputes the matrix $M=2AX$, where $M_{i,m}$ is the total similarity of object $i$ to cluster $m$. Precomputing $M$ takes $O(kn^{2})$, but allows gradient computation in $O(k)$ (line 8). We then have to update $M_{i,m}$ accordingly (lines 15 and 18), which is $O(n)$, reducing the per-iteration complexity to $O(n+k)$. Thus, the total complexity is $O(kn^{2}+T(n+k))$, which improves on the $O(Tnk)$ approach because, (i) computing $M$ involves a sparse matrix product, which is highly efficient in practice, and (ii) since $T>n$, reducing per-iteration cost leads to significant practical gains.
3.5 Impact of $\alpha$ and $\beta$
We now analyze the impact of $\alpha$ and $\beta$ in Eq. 7. We begin by stating the following proposition.
Proposition 4.
(a) There exists a $\xi_{1}<0$ such that for any $\beta\leq\xi_{1}$, there is a clustering solution maximizing Eq. 7 where all the objects are assigned to a single non-neutral cluster. (b) Conversely, there exists a $\xi_{2}>0$ such that for any $\beta\geq\xi_{2}$, there is a clustering solution maximizing Eq. 7 where all the objects are neutral.
From Proposition 4, we understand the extreme cases of $\beta$: (a) a small negative $\beta$ results in a maximally imbalanced non-neutral clustering (i.e., all objects in one non-neutral cluster), while (b) a large positive $\beta$ makes all objects neutral. For intermediate $\beta\in[\xi_{1},\xi_{2}]$, we analyze the gradient in Eq. 13. Increasing $\beta$ strictly reduces the contribution of object $i$ to each cluster $m\in[k]$, but since the term $-2\beta|S_{m}|$ scales with cluster size, larger clusters become less favorable, promoting balance. If $\beta$ is large enough, it forces $G_{i,m}<0$ for all $m\in[k]$, making neutrality optimal for object $i$. Note that this is more likely for low-degree objects, implying that high-degree objects (with clear cluster assignment) are more likely to remain non-neutral, resulting in dense non-neutral clusters. Consequently, increasing $\beta$ leads to smaller (i.e., more neutral objects) and denser non-neutral clusters, while maintaining balanced, as desired.
The parameter $\alpha$ has been studied in prior work (Chu et al., 2016; Tzeng et al., 2020). From Eq. 7, $\alpha$ balances maximizing intra-similarities and minimizing inter-similarities, which translates to a trade-off between cohesion within clusters and separation between them. A heuristic choice of $\alpha=1/(k-1)$ was proposed in (Tzeng et al., 2020), based on the observation that the number of intra-similarities scale linearly with $k$, while the number of inter-similarities grow quadratically. This choice prevents inter-similarities from dominating the objective. Finally, the term $-\alpha\eta_{i}$ indicates that $\alpha$ influences whether object $i$ becomes neutral, underscoring the need to account for inter-similarities in the objective (as suggested in Section 2.3).
4 Experiments
Table 1: Results for different methods and real-world datasets w.r.t. balance-aware polarity. $|E|$ is number of non-zero edges. Dashes indicate the method exceeded memory capacity. LSPCD (ours) yields the best results in most of the cases.
W8 BTC WikiV REF SD WikiC EP WikiP
$|V|$ 790 5.9K 7.1K 10.9K 82.1K 116.7K 131.6K 138.6K
$|E|$ 116K 214.5K 1M 251.4K 500.5K 2M 711.2K 715.9K
$k=4$ LSPCD (ours) 6.9 3.6 5.9 3.9 6.7 26.3 14.9 5.4
SCG-MA 5.8 0.3 0.8 0.3 1.9 0.7 0.1 0.7
SCG-MO 4.9 0.3 0.4 0.4 0.2 6.6 0.1 0.0
SCG-B 0.3 0.1 0.2 0.1 0.5 18.5 0.2 0.1
SCG-R 1.3 0.9 0.9 0.6 2.4 2.0 0.3 1.0
KOCG-top-$1$ 5.8 4.2 1.2 2.6 0.9 0.4 4.0 0.9
KOCG-top-$r$ 5.4 2.7 2.6 1.6 1.0 2.8 5.9 1.2
BNC-$(k+1)$0.1 -0.8 -0.4 -1.0
BNC-$k$ 0.5 0.0 0.0 0.0
SPONGE-$(k+1)$ 22.9 0.0 0.4 0.3
SPONGE-$k$ 18.6 0.0 0.0 0.0
$k=6$ LSPCD (ours) 4.4 2.2 3.6 2.2 3.8 16.4 6.8 3.5
SCG-MA 3.1 0.1 0.2 0.1 0.0 0.6 0.4 0.0
SCG-MO 3.6 0.1 0.2 0.1 0.1 4.2 0.1 0.0
SCG-B 0.3 0.2 0.2 0.1 0.2 1.8 0.2 0.1
SCG-R 0.3 0.3 0.4 0.2 0.4 2.3 0.1 0.0
KOCG-top-$1$ 6.1 1.7 2.5 3.3 1.7 0.4 2.7 3.7
KOCG-top-$r$ 5.0 1.2 1.8 2.5 1.3 1.1 2.8 0.7
BNC-$(k+1)$0.1 -0.1 -0.8 -0.4
BNC-$k$ 0.5 0.0 0.0 0.0
SPONGE-$(k+1)$ 17.7 0.0 0.4 0.6
SPONGE-$k$ 16.0 0.0 0.0 0.0
Table 2: Results for different methods and real-world datasets w.r.t. polarity. LSPCD (ours) yields the best or close to the best results. $|E|$ is number of non-zero edges. Dashes indicate the method exceeded memory capacity. LSPCD (ours) yields the best results in most of the cases.
W8 BTC WikiV REF SD WikiC EP WikiP
$|V|$ 790 5.9K 7.1K 10.9K 82.1K 116.7K 131.6K 138.6K
$|E|$ 116K 214.5K 1M 251.4K 500.5K 2M 711.2K 715.9K
$k=4$ LSPCD (ours) 218.5 23.3 52.6 139.2 61.1 113.6 111.5 71.6
SCG-MA 205.1 25.1 52.9 94.5 35.5 104.9 127.4 56.5
SCG-MO 213.2 25.3 53.1 82.1 38.5 117.9 129.0 39.7
SCG-B 211.6 12.4 24.8 116.2 48.3 49.8 94.4 45.7
SCG-R 214.6 8.0 19.5 118.7 10.7 41.1 65.1 33.7
KOCG-top-$1$ 9.1 8.4 4.5 15.0 2.6 4.5 8.9 3.1
KOCG-top-$r$ 7.4 5.0 3.3 3.7 3.0 3.8 11.0 4.4
BNC-$(k+1)$0.2 -9.4 -1.1 -1.0
BNC-$k$ 185.3 5.2 15.8 41.5
SPONGE-$(k+1)$ 53.8 1.1 1.0 1.0
SPONGE-$k$ 71.2 5.1 15.8 41.5
$k=6$ LSPCD (ours) 217.3 20.0 46.2 137.6 57.1 96.1 103.4 58.7
SCG-MA 207.3 14.6 45.5 84.9 37.8 102.6 88.8 57.5
SCG-MO 205.8 15.2 47.0 55.6 34.6 111.6 129.2 41.8
SCG-B 211.6 9.3 23.3 116.2 47.7 46.1 94.5 46.0
SCG-R 201.2 6.9 10.4 50.3 7.9 18.3 43.3 3.3
KOCG-top-$1$ 7.9 4.1 4.5 8.6 3.6 4.9 6.0 10.1
KOCG-top-$r$ 9.1 3.6 3.1 4.0 3.3 1.5 6.8 3.6
BNC-$(k+1)$0.2 -4.2 -1.1 -0.8
BNC-$k$ 185.2 5.2 15.8 41.5
SPONGE-$(k+1)$ 47.8 1.3 1.0 1.0
SPONGE-$k$ 57.9 5.1 15.8 41.5
In this section, we describe our experimental studies, where extensive additional results are presented in Appendix B. We use eight publically available datasets, which are commonly used in previous work for PCD. In Appendix B.1, we provide details about each dataset. We compare our efficient local search method for PCD, called LSPCD, against a number of baselines introduced below.
Baselines. (i) SCG (Tzeng et al., 2020) is a spectral method that identifies $k$ non-neutral clusters by maximizing polarity (Eq. 6) with $\alpha=1/(k-1)$. It solves a continuous relaxation and applies one of four rounding techniques, resulting in SCG-MA, SCG-R, SCG-MO, and SCG-B. We refer to (Tzeng et al., 2020) for details. (ii) KOCG (Chu et al., 2016) optimizes a similar objective and formulates it as a constrained quadratic optimization problem (this optimization approach is very different from ours). It outputs a set of local minima. For comparison, we select KOCG-top-$1$ (the best local minimum) and KOCG-top-$r$, where $r$ is chosen such that the number of non-neutral objects is closest to SCG-MA, following (Tzeng et al., 2020). (iii) BNC (Chiang et al., 2012) and SPONGE (Cucuringu et al., 2019) are spectral methods designed for SNP that do not explicitly handle neutral objects. As in (Tzeng et al., 2020), we apply two heuristics with these methods: (a) we treat all $k$ clusters as non-neutral, and (b) we run the methods with $k+1$ clusters and then designate the largest cluster as neutral. These variants are denoted BNC-$k$ / SPONGE-$k$ and BNC-$(k+1)$ / SPONGE-$(k+1)$, respectively. We use publicly available implementations of these baselines, with further details, including hyperparameters, provided in Appendix B.2.
Metrics. (i) Following prior work, we use polarity to evaluate the quality of different methods (Bonchi et al., 2019; Tzeng et al., 2020), defined as in Eq. 6 with $\alpha=1/(k-1)$. We note that SCG (Tzeng et al., 2020) explicitly optimizes for polarity, unlike our approach, which gives it an inherent advantage in polarity-based comparisons. (ii) We introduce balance-aware polarity (BA-Polarity), which multiplies polarity by a balance factor $|S_{m}|/|S_{l}|$, where $|S_{m}|$ and $|S_{l}|$ are the sizes of the smallest and largest non-neutral clusters, respectively. In Appendix B.3, we provide a detailed presentation of the solutions found by each method. This includes the number of non-neutral objects identified, the balance of the non-neutral clusters, the runtime, and more.
Results. Tables 1 and 2 present results for different methods and datasets with $k=4$ and $k=6$, evaluating BA-Polarity and polarity, respectively. The spectral clustering methods, BNC and SPONGE, exceeded memory limits on large datasets (caused by $k$-means), indicated by dashes. Unlike most prior work, which primarily focuses on $k=2$, we consider more than two clusters; however, for completeness, results for $k=2$ are provided in Appendix B.3, where similar trends are observed. We report the mean over five runs with different seeds, with standard deviations included in Appendix B.3. For our method, we select the $\beta$ value that maximizes polarity, testing 10 values per dataset, while we fix $\alpha=1/(k-1)$ for all methods unless stated otherwise. Results show that our method is highly competitive—often the best—in polarity across all datasets, despite SCG explicitly optimizing for polarity while our method does not. Moreover, our method consistently outperforms all baselines in BA-Polarity. While SCG remains competitive in polarity, our method surpasses it in BA-Polarity across all datasets. In contrast, KOCG and SPONGE occasionally achieve higher BA-Polarity but at the cost of extremely low polarity (see Table 2) due to overly balanced clustering solutions (arguably to an excessive degree). This demonstrates that our method finds high-polarity solutions while maintaining balance, effectively addressing SCG’s tendency to produce highly imbalanced clusters. Additionally, our method does not impose strict balance constraints, which is beneficial since real-world clusterings are rarely perfectly balanced. Instead, it identifies high-polarity solutions with reasonable balance, making it more practical for real-world applications (further discussed in Appendix B.3).
Figure 1: Impact of $\beta$ on the WikiPol dataset with $k=6$.
Figure 1 shows the effect of varying $\beta$. Very small or large $\beta$ values lead to poorer polarity, as they produce clustering solutions with too many or too few of non-neutral objects, respectively. In contrast, intermediate $\beta$ values yield competitive polarity while consistently outperforming baselines in BA-Polarity. Further insights on the impact of $\beta$ and $\alpha$ are provided in Appendix B.3, along with other results.
5 Conclusion
We proposed a novel formulation for the polarized community discovery problem, emphasizing (i) large, (ii) dense, and (iii) balanced clustering solutions. We developed an efficient and effective local search method and established its connection to block-coordinate Frank-Wolfe optimization, proving a linear convergence rate of $O(1/t)$. Our extensive experimental results demonstrate that our method achieves high-polarity clustering solutions while maintaining balance, and outperforms the other methods.
Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A Proofs
See 1
Proof.
We begin by defining the following quantities, which are constants w.r.t. different clustering solutions for the $k$-CC problem.
$c_{\text{sim}}\triangleq\sum_{(i,j)\in E}A_{i,j}.$ (14)
$c_{\text{abs}}\triangleq\sum_{(i,j)\in E}|A_{i,j}|.$ (15)
The five objectives can be written as follows.
$f^{\text{full}}$ $\triangleq N^{+}_{\text{intra}}-N^{-}_{\text{intra}}+N^{-}_{\text{inter}}-N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$ (16)
$f^{\text{MaxAgree}}$ $\triangleq N^{+}_{\text{intra}}+N^{-}_{\text{inter}}$
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{+}_{i,j}-\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A^{-}_{i,j}$
$=\frac{1}{2}\sum_{m\in[k]}\sum_{i,j\in S_{m}}(|A_{i,j}|+A_{i,j})-\frac{1}{2}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}(|A_{i,j}|-A_{i,j})$
$f^{\text{MinDisagree}}$ $\triangleq N^{-}_{\text{intra}}+N^{+}_{\text{inter}}$
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{-}_{i,j}+\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A^{+}_{i,j}$
$=\frac{1}{2}\sum_{m\in[k]}\sum_{i,j\in S_{m}}(|A_{i,j}|-A_{i,j})+\frac{1}{2}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}(|A_{i,j}|+A_{i,j})$
$f^{\text{MaxCorr}}$ $\triangleq N^{+}_{\text{intra}}-N^{-}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}$
$f^{\text{MinCut}}$ $\triangleq-N^{-}_{\text{inter}}+N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$
Given this, we observe the following connection between the objectives.
$f^{\text{MaxAgree}}$ $=\frac{1}{2}\sum_{m\in[k]}\sum_{i,j\in S_{m}}(|A_{i,j}|+A_{i,j})-\frac{1}{2}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}(|A_{i,j}|-A_{i,j})$ (17)
$=\frac{1}{2}f^{\text{full}}+\frac{1}{2}c_{\text{abs}}$
$=-f^{\text{MinDisagree}}+c_{\text{abs}}$
$=\frac{1}{2}f^{\text{MaxCorr}}-\frac{1}{2}f^{\text{MinCut}}+\frac{1}{2}c_{\text{abs}}$
$=f^{\text{MaxCorr}}-\frac{1}{2}c_{\text{sim}}+\frac{1}{2}c_{\text{abs}}$
$=-f^{\text{MinCut}}+\frac{1}{2}c_{\text{sim}}+\frac{1}{2}c_{\text{abs}}$
The above establishes that they are all equal up to constants. We prove the last statement of the proposition by counterexample. We consider the graph $V=\{1,2,3\}$ with edge weights $A_{1,2}=+1$, $A_{2,3}=+1$, $A_{1,3}=-1$. The possible clustering solutions (partitions) are:
$S^{(1)}=\{\{1,2,3\}\},\quad S^{(2)}=\{\{1,2\},\,\{3\}\},\quad S^{(3)}=\{\{1,3\},\,\{2\}\},\quad S^{(4)}=\{\{2,3\},\,\{1\}\},\quad S^{(5)}=\{\{1\},\,\{2\},\,\{3\}\}.$
In Table 3, we list all linear combinations of the terms $N_{\text{intra}}^{+},-N_{\text{intra}}^{-},N_{\text{inter}}^{-},-N_{\text{inter}}^{+}$ evaluated on each of the five clustering solutions. Expectedly (from the first part of the proposition), the five objectives $f^{\text{full}},f^{\text{MaxAgree}},f^{\text{MinDisagree}},f^{\text{MaxCorr}},f^{\text{MinCut}}$ all produce the same ranking of these solutions. In contrast, every other combination of the terms ranks at least one solution differently compared to these five. This proves the last statement of the proposition.
Table 3: All sums of $N_{\text{intra}}^{+},\,-N_{\text{intra}}^{-},\,N_{\text{inter}}^{-},\,-N_{\text{inter}}^{+}$ and their values on the five partitions $S^{(m)}$. In parentheses we indicate the known name when the combination corresponds to one of the five standard correlation-clustering objectives (or its negative).
Idx Combination $S^{(1)}$ $S^{(2)}$ $S^{(3)}$ $S^{(4)}$ $S^{(5)}$
1 $N_{\text{intra}}^{+}$ 2 1 0 1 0
2 $-\,N_{\text{intra}}^{-}$1 0 -1 0 0
3 $N_{\text{inter}}^{-}$ 0 1 0 1 1
4 $-\,N_{\text{inter}}^{+}$ 0 -1 -2 -1 -2
5 $N_{\text{intra}}^{+}-N_{\text{intra}}^{-}$ ($f^{\text{MaxCorr}}$) 1 1 -1 1 0
6 $N_{\text{intra}}^{+}+N_{\text{inter}}^{-}$ ($f^{\text{MaxAgree}}$) 2 2 0 2 1
7 $N_{\text{intra}}^{+}-N_{\text{inter}}^{+}$ 2 0 -2 0 -2
8 $-\,N_{\text{intra}}^{-}+N_{\text{inter}}^{-}$1 1 -1 1 1
9 $-\,N_{\text{intra}}^{-}-N_{\text{inter}}^{+}$ ($-\,f^{\text{MinDisagree}}$) -1 -1 -3 -1 -2
10 $N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$ ($-f^{\text{MinCut}}$) 0 0 -2 0 -1
11 $N_{\text{intra}}^{+}-N_{\text{intra}}^{-}+N_{\text{inter}}^{-}$ 1 2 -1 2 1
12 $N_{\text{intra}}^{+}-N_{\text{intra}}^{-}-N_{\text{inter}}^{+}$ 1 0 -3 0 -2
13 $N_{\text{intra}}^{+}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$ 2 1 -2 1 -1
14 $-\,N_{\text{intra}}^{-}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$1 0 -3 0 -1
15 $N_{\text{intra}}^{+}-N_{\text{intra}}^{-}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$ ($f^{\text{full}}$) 1 1 -3 1 -1
See 2
Proof.
If an object transitions from neutral to non-neutral, it may introduce agreements (positive intra-cluster or negative inter-cluster similarities) and/or disagreements (negative intra-cluster or positive inter-cluster similarities). An exception is objects with zero degree (zero similarity to all others), which can be assigned as neutral or non-neutral without affecting any of the five objectives. Thus, we only consider non-zero degree objects in the remainder of the proof.
The max agreement objective (Eq. 2) considers only agreements. Making an object non-neutral either increases or maintains the objective but never decreases it, ensuring all objects become non-neutral. Conversely, the min disagreement objective (Eq. 3) considers only disagreements. Making an object non-neutral either decreases or maintains the objective but never improves it, ensuring all objects remain neutral.
Now, consider a clustering with $k$ non-neutral clusters, where all intra-cluster similarities are $+1$ and all inter-cluster similarities are $-1$. If an unassigned object $i\in V$ has similarity $+1$ to all others, the max correlation objective (Eq. 4) assigns it to the largest non-neutral cluster, while the minimum cut objective (Eq. 5) keeps it neutral. For $k>2$, the full objective (Eq. 1) places $i$ in the largest non-neutral cluster if its size exceeds the sum of all others; otherwise, it remains neutral. Conversely, if object $i$ has similarity $-1$ to all others, max correlation keeps it neutral, whereas minimum cut assigns it to the smallest non-neutral cluster. For $k>2$, the full objective may assign $i$ as neutral or non-neutral depending on cluster sizes.
Therefore, we conclude that max agreement and min disagreement differ fundamentally, always assigning all objects as non-neutral or neutral, respectively. Furthermore, from the two counterexamples above, max correlation, minimum cut, and the full objective are not equivalent and none of them guarantee that all objects are either neutral or non-neutral in all cases (meaning they are all different from max agreement and min disagreement in general).
For $k=2$, the full objective always increases (or remains constant) when an object is made non-neutral, aligning it with max agreement. To see this, consider a clustering with $k=2$ non-neutral clusters, and let $M_{i,m}=\sum_{j\in S_{m}}A_{i,j}$ be the total similarity of object $i$ to cluster $m$. The impact on the objective when assigning $i$ to $m$ is $M_{i,m}-M_{i,p}$, where $p$ is the other cluster. Since this difference is always positive when $i$ is placed in its most similar cluster, assigning $i$ as non-neutral always improves the objective. Then, since all objects are non-neutral, the problem is equivalent to the $k$-CC problem where we know max agreement and the full objective are equivalent (from Proposition 1). For $k>2$, this reasoning no longer holds, as contributions from other clusters can outweigh the within-cluster similarity to the most similar cluster (i.e., making the total contribution negative), potentially making neutrality optimal. However, we note that in our final objective (Eq. 7), when $\alpha$ and $\beta$ are involved, all terms will contribute with unique information even for $k=2$.
See 1
Proof.
Fix $\alpha,\beta\in\mathbb{R}$ to any values. Assume that we know which objects in $V$ should be assigned to the neutral set $S_{0}$ in the optimal solution to the $k$-PCD problem. Then, let $V^{\prime}=V\setminus S_{0}$ and let $E^{\prime}$ be the set of edges between objects in $V^{\prime}$. Since no object in $V^{\prime}$ should be neutral, the problem reduces to finding a partition of $V^{\prime}$ that maximizes Eq. 7. We rewrite our objective in Eq. 7 as
$(N^{+}_{\text{intra}}-N^{-}_{\text{intra}})+\alpha(N^{-}_{\text{inter}}-N^{+}_{\text{inter}})-\beta\sum_{m\in[k]}|S_{m}|^{2}$ $=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}-\beta\sum_{m\in[k]}|S_{m}|^{2}$ (18)
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}.$
The second equality follows from Proposition 3. Defining $c_{\text{sim}}\triangleq\sum_{(i,j)\in E^{\prime}}A_{i,j}$, we obtain:
$\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}=c_{\text{sim}}-\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}.$ (19)
Substituting this into Eq. 18 and simplifying:
$\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$ $=\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha(c_{\text{sim}}-\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j})$ (20)
$=(1+\alpha)\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha c_{\text{sim}}.$
Defining $A^{\prime}_{i,j}=(1+\alpha)(A_{i,j}-\beta)$, we observe that since $c_{\text{sim}}$ is a constant across clustering solutions, the problem reduces to finding a partition of $V^{\prime}$ that maximizes $\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{\prime}_{i,j}$. This is equivalent to the max correlation objective (Eq. 4) applied to the transformed adjacency matrix $A^{\prime}$. By Proposition 1, this objective is equivalent to the $k$-CC problem (Problem 1). Thus, solving the $k$-PCD problem requires solving the $k$-CC problem on the instance $G^{\prime}=(V^{\prime},E^{\prime})$, meaning $k$-PCD is at least as hard as $k$-CC. Since correlation clustering is NP-hard [Bansal et al., 2004, Giotis and Guruswami, 2006], we conclude that $k$-PCD is also NP-hard. ∎
See 3
Proof.
We have
$(N^{+}_{\text{intra}}-N^{-}_{\text{intra}})+\alpha(N^{-}_{\text{inter}}-N^{+}_{\text{inter}})-\beta\sum_{m\in[k]}|S_{m}|^{2}=$ (21)
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}-\beta\sum_{m\in[k]}|S_{m}|^{2}$
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}-\beta\sum_{m\in[k]}\sum_{i,j\in S_{m}}1$
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}-\sum_{m\in[k]}\sum_{i,j\in S_{m}}\beta$
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}.$
The second equality (line 3) holds because the number of pairs of objects inside cluster $m$ is $|S_{m}|^{2}$. A similar regularization is established in [Chehreghani, 2023] for the minimum cut objective, where it is shown that optimizing this minimum cut objective regularized with $-\beta\sum_{m\in[k]}|S_{m}|^{2}$ is equivalent to optimizing the max correlation objective (Eq. 4) with similarities shifted by $\beta$. However, their result specifically considers the full network partitioning of unsigned networks, where the initial pairwise similarities are assumed non-negative. Moreover, they use a different regularization in practice: they shift the pairwise similarities so that the sum of the rows and columns of the similarity matrix becomes zero. ∎
See 2
Proof.
We begin by writing our objective function $f({\bm{x}}_{[n]})$ in Eq. 10 as follows.
$\begin{split}f({\bm{x}}_{[n]})&=\sum_{(i,j)\in E}\sum_{m\in[k]}x_{im}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{(i,j)\in E}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}x_{im}x_{jp}A_{i,j}\\ &=-\sum_{i\in[n]}\sum_{m\in[k]}x^{2}_{im}\beta+\sum_{\begin{subarray}{c}(i,j)\in E\\ i\neq j\end{subarray}}\sum_{m\in[k]}x_{im}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{(i,j)\in E}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}x_{im}x_{jp}A_{i,j}\\ &=-\sum_{i\in[n]}\sum_{m\in[k]}x_{im}\beta+\sum_{\begin{subarray}{c}(i,j)\in E\\ i\neq j\end{subarray}}\sum_{m\in[k]}x_{im}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{(i,j)\in E}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}x_{im}x_{jp}A_{i,j}.\end{split}$ (22)
In the second equality, we separate out the terms for $i=j$ and use that $A_{i,i}=0$. In the third equality, we consider that ${\bm{x}}_{[n]}$ is a discrete solution. This makes the first term linear instead of being quadratic w.r.t. $x_{im}$, which is a crucial step in proving the theorem. Let $f({\bm{x}}_{i})$ denote $f({\bm{x}}_{[n]})$ when treating all blocks other than ${\bm{x}}_{i}$ as constants. Then,
$\begin{split}f({\bm{x}}_{i})&=-\sum_{m\in[k]}x_{im}\beta+2\sum_{j\in[n]\setminus\{i\}}\sum_{m\in[k]}x_{im}x_{jm}(A_{i,j}-\beta)-2\alpha\sum_{j\in[n]\setminus\{i\}}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}x_{im}x_{jp}A_{i,j}+C\\ &=\sum_{m\in[k]}x_{im}\underbrace{\Big{(}-\beta+2\sum_{j\in[n]\setminus\{i\}}\big{(}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{p\in[k]\setminus\{m\}}x_{jp}A_{i,j}\big{)}\Big{)}}_{c_{im}}+C,\end{split}$ (23)
where $C$ denotes terms independent of ${\bm{x}}_{i}$. Define ${\bm{c}}_{i}\in\mathbb{R}^{k+1}$ with elements
$c_{im}\triangleq-\beta+2\sum_{j\in[n]\setminus\{i\}}\big{(}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{p\in[k]\setminus\{m\}}x_{jp}A_{i,j}\big{)},\quad\text{for }m\in[k],\quad c_{i0}\triangleq 0.$ (24)
Then, we obtain
$f({\bm{x}}_{i})=\sum_{m\in\{0,\dots,k\}}x_{im}c_{im}+C={\bm{x}}_{i}^{T}{\bm{c}}_{i}+C.$ (25)
Eq. 25 clearly illustrates that the contribution of the neutral component (index zero) of each $x_{im}$ is not included in the total objective (since $c_{i0}=0$). From Eq. 25, the gradient of $f({\bm{x}}_{[n]})$ w.r.t. ${\bm{x}}_{i}$ is
$\nabla_{i}f({\bm{x}}_{[n]})={\bm{c}}_{i}.$ (26)
Let ${\bm{c}}^{(t)}_{i}=\nabla_{i}f({\bm{x}}^{(t)}_{[n]})$ be the gradient of $f({\bm{x}}_{[n]})$ evaluated at the current solution ${\bm{x}}_{[n]}^{(t)}$ (defined as in Eq. 24). The optimization problem on line 4 of Algorithm 1 is
${\bm{x}}_{i}^{\ast}=\operatorname*{arg\,max}_{{\bm{x}}_{i}\in\Delta^{k+1}}{\bm{x}}_{i}^{T}{\bm{c}}_{i}^{(t)}.$ (27)
Since Eq. 27 is a linear program over the simplex $\Delta^{k+1}$, the optimal solution is obtained by setting $x_{im}^{\ast}=1$ for $m=\operatorname*{arg\,max}_{m\in\{0,\dots,k\}}c^{(t)}_{im}$ and $x_{ip}^{\ast}=0$ for all $p\neq m$. This proves the first statement of part (a) of the theorem.
Next, we note that the difference $f({\bm{x}}^{\ast}_{[n]})-f({\bm{x}}^{(t)}_{[n]})$ simplifies to $f({\bm{x}}^{\ast}_{i})-f({\bm{x}}^{(t)}_{i})$ (where $f({\bm{x}}_{i})$ is defined in Eq. 23), since only the terms involving the variables in block $i$ change between ${\bm{x}}^{\ast}_{[n]}$ and ${\bm{x}}^{(t)}_{[n]}$. Therefore, we can derive the following.
$f({\bm{x}}^{\ast}_{[n]})-f({\bm{x}}^{(t)}_{[n]})$ $=f({\bm{x}}^{\ast}_{i})-f({\bm{x}}^{(t)}_{i})$ (28)
$=(({\bm{x}}_{i}^{\ast})^{T}{\bm{c}}_{i}^{\ast}+C)-(({\bm{x}}_{i}^{(t)})^{T}{\bm{c}}_{i}^{(t)}+C)$
$=(({\bm{x}}_{i}^{\ast})^{T}{\bm{c}}_{i}^{(t)}+C)-(({\bm{x}}_{i}^{(t)})^{T}{\bm{c}}_{i}^{(t)}+C)$
$=(({\bm{x}}_{i}^{\ast})^{T}-({\bm{x}}_{i}^{(t)})^{T}){\bm{c}}_{i}^{(t)}$
$=({\bm{x}}_{i}^{\ast}-{\bm{x}}_{i}^{(t)})\cdot\nabla_{i}f({\bm{x}}^{(t)}_{[n]})$
Here, ${\bm{c}}_{i}^{\ast}$ is defined as in Eq. 24 w.r.t. ${\bm{x}}^{\ast}_{[n]}$. Since ${\bm{x}}^{\ast}_{[n]}$ and ${\bm{x}}^{(t)}_{[n]}$ differ only in block $i$, and neither ${\bm{c}}_{i}^{\ast}$ nor ${\bm{c}}_{i}^{(t)}$ depend on the variables in block $i$, it follows that ${\bm{c}}_{i}^{\ast}={\bm{c}}_{i}^{(t)}$, justifying the third equality. In Eq. 22, we assume that ${\bm{x}}_{[n]}$ is discrete. To ensure this property holds throughout, we require that both ${\bm{x}}^{\ast}_{[n]}$ and ${\bm{x}}_{[n]}^{(t)}$ remain discrete for all $t\in\{0,\dots,T\}$.
First, by assumption in the theorem, ${\bm{x}}_{[n]}^{(0)}$ is discrete. From part (a), we know that ${\bm{x}}_{i}^{\ast}$ is discrete, implying ${\bm{x}}^{\ast}_{[n]}$ is discrete as long as ${\bm{x}}_{[n]}^{(t)}$ is discrete. Furthermore, from Eq. 28, the optimal solution ${\bm{x}}_{i}^{\ast}$ in line 4 of Algorithm 1 maximally increases the objective, which ensures the optimal step size in line 6 is $\gamma=1$ (proving the second statement of part (a)). Consequently, ${\bm{x}}_{i}^{(t+1)}$ remains discrete. By induction, this guarantees that ${\bm{x}}^{(t)}_{[n]}$ is discrete for all $t$, ensuring Eq. 28 holds for all $t\in\{0,\dots,T\}$. This completes the proof of part (b) of the theorem.
See 3
Proof.
From Definition 1, we have that, in our case, the FW duality gap is defined as
$g({\bm{x}}_{[n]})\triangleq\max_{{\bm{s}}_{i}\in\Delta^{k+1},\forall i\in[n]}({\bm{s}}_{[n]}-{\bm{x}}_{[n]})\cdot\nabla f({\bm{x}}_{[n]}).$ (29)
Then, we recall that
$\tilde{g}_{t}=\min_{0\leq l\leq t-1}g({\bm{x}}^{(l)}_{[n]})$ (30)
is the smallest duality gap observed in Alg. 1 up until step $t$. As established by [Lacoste-Julien et al., 2013] for general domains, the FW duality gap can be decomposed as follows.
$g({\bm{x}}_{[n]})$ $\triangleq\max_{{\bm{s}}_{i}\in\Delta^{k+1},\forall i\in[n]}({\bm{s}}_{[n]}-{\bm{x}}_{[n]})\cdot\nabla f({\bm{x}}_{[n]})$ (31)
$=\max_{{\bm{s}}_{i}\in\Delta^{k+1},\forall i\in[n]}\sum_{i\in[n]}({\bm{s}}_{i}-{\bm{x}}_{i})\cdot\nabla_{i}f({\bm{x}}_{[n]})$
$=\sum_{i\in[n]}\underbrace{\max_{{\bm{s}}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}_{i})\cdot\nabla_{i}f({\bm{x}}_{[n]})}_{\triangleq g_{i}({\bm{x}}_{[n]})}$
Let $g_{i}({\bm{x}}_{[n]})\triangleq\max_{{\bm{s}}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}_{i})\cdot\nabla_{i}f({\bm{x}}_{[n]})$ be the duality gap related to block $i$. We have that the FW duality gap is the sum of the gaps from each block: $g({\bm{x}}_{[n]})=\sum_{i\in[n]}g_{i}({\bm{x}}_{[n]})$.
From Definition 2 (convergence rate), in order to prove the stated convergence rate, we need to show that $\mathbb{E}[\tilde{g}_{t}]\leq nh_{0}/t$. The structure of our proof is similar to the proof of Theorem 2 in [Thiel et al., 2019]. However, here we adapt it to our problem and make the proof more rigorous (including correction of a mistake in the proof by [Thiel et al., 2019]). A key difference is that our objective in Eq. 10 is not multilinear in the blocks $i$. Then, as shown in Eq. 22 of Theorem 2, the first quadratic term can be transformed into a linear one by assuming a discrete solution (which we showed holds at every step $t$).
In Alg. 1, a block $i\in[n]$ is chosen uniformly at random (on line 3). Therefore, we have
$\mathbb{E}[g_{i}({\bm{x}}^{(t)}_{[n]})|{\bm{x}}^{(t)}_{[n]}]$ $=\sum_{i\in[n]}P(i\text{ is selected})g_{i}({\bm{x}}^{(t)}_{[n]})$ (32)
$=\sum_{i\in[n]}\frac{1}{n}g_{i}({\bm{x}}^{(t)}_{[n]})$
$=\frac{1}{n}\sum_{i\in[n]}\max_{{\bm{s}}^{(t)}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}^{(t)}_{i})\cdot\nabla_{i}f({\bm{x}}^{(t)}_{[n]})$
$=\frac{1}{n}g({\bm{x}}_{[n]}^{(t)}).$
We now take an expectation w.r.t. ${\bm{x}}^{(t)}_{[n]}$ on both sides and obtain
[…]
$=\mathbb{E}[g_{i}({\bm{x}}_{[n]}^{(t)})],$
where the last equality follows from the Law of Total Expectation (i.e., that $\mathbb{E}_{Y}[\mathbb{E}_{X}{[X|Y]}]=\mathbb{E}_{X}{[X}]$, where $X$ and $Y$ are random variables). We therefore have that $\frac{1}{n}\mathbb{E}[g({\bm{x}}_{[n]}^{(t)})]=\mathbb{E}[g_{i}({\bm{x}}_{[n]}^{(t)})]$, where the expectation is w.r.t. all randomly chosen blocks $i$ before step $t$. Now, from Theorem 2 we have that our objective satisfies
$f({\bm{x}}^{(t+1)}_{[n]})-f({\bm{x}}^{(t)}_{[n]})=g_{i}({\bm{x}}_{[n]}^{(t)})=\max_{{\bm{s}}^{(t)}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}^{(t)}_{i})\cdot\nabla_{i}f({\bm{x}}^{(t)}_{[n]}).$ (34)
Then, we have
$\frac{1}{n}\sum_{t=0}^{T-1}\mathbb{E}[g({\bm{x}}^{(t)})]$ $=\sum_{t=0}^{T-1}\mathbb{E}[g_{i}({\bm{x}}_{[n]}^{(t)})]$ (35)
$=\sum_{t=0}^{T-1}\mathbb{E}[f({\bm{x}}^{(t+1)}_{[n]})-f({\bm{x}}^{(t)}_{[n]})]$
$=\mathbb{E}[f({\bm{x}}^{(T)}_{[n]})]-f({\bm{x}}^{(0)}_{[n]})$
$\leq OPT-f({\bm{x}}_{[n]}^{(0)}),$
where the third equality is due to the telescoping rule and $OPT$ is the objective value of the optimal clustering solution to Problem 1 ($k$-PCD). On the other hand, we have
$\frac{1}{n}\sum_{t=0}^{T-1}\mathbb{E}[g({\bm{x}}^{(t)})]\geq\frac{T}{n}\mathbb{E}[\tilde{g}_{T}],$ (36)
where $\tilde{g}_{t}$ is defined as in Eq. 30 (the smallest gap observed until step $t$). Therefore,
$\frac{T}{n}\mathbb{E}[\tilde{g}_{T}]$ $\leq OPT-f({\bm{x}}^{(0)})$ (37)
$\Rightarrow\mathbb{E}[\tilde{g}_{T}]$ $\leq\frac{n(OPT-f({\bm{x}}^{(0)}))}{T}.$
The value of $OPT$ depends on the particular instance. In order to obtain an instance-independent bound, we use that $OPT-f({\bm{x}}^{(0)})\leq\sum_{(i,j)\in E}|A_{i,j}|$ resulting in
$\mathbb{E}[\tilde{g}_{T}]\leq\frac{n\sum_{(i,j)\in E}|A_{i,j}|}{T}.$ (38)
which we aimed to show since it holds for any $T$. ∎
See 1
Proof.
From part (a) of Theorem 2, the current solution, ${\bm{x}}^{(t)}_{[n]}$, remains discrete (i.e., hard cluster assignments) at every step of Alg. 1 for all $i\in[n]$. Moreover, each step of Alg. 1 consists of placing object $i$ in the cluster $m\in\{0,\dots,k\}$ with maximal gradient $G_{i,m}$. By part (b) of Theorem 2, this is equivalent to placing object $i$ in the cluster that maximally improves our objective in Eq. 7. ∎
See 4
Proof.
From Theorem 2, we recall that since the current solution ${\bm{x}}_{[n]}^{(t)}$ always remains discrete, our objective can be written as
$f({\bm{x}}^{(t)}_{[n]})=-\sum_{i\in[n]}\sum_{m\in[k]}x^{(t)}_{im}\beta+\sum_{\begin{subarray}{c}(i,j)\in E\\ i\neq j\end{subarray}}\sum_{m\in[k]}x^{(t)}_{im}x^{(t)}_{jm}(A_{i,j}-\beta)-\alpha\sum_{(i,j)\in E}\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}x^{(t)}_{im}x^{(t)}_{jp}A_{i,j}.$ (39)
We let $f({\bm{x}}^{(t)}_{i})$ denote $f({\bm{x}}^{(t)}_{[n]})$ when treating all blocks other than ${\bm{x}}^{(t)}_{i}$ as constants. Then,
$f({\bm{x}}^{(t)}_{i})=\sum_{m\in[k]}x^{(t)}_{im}\big{(}\underbrace{-\beta+2\sum_{j\in[n]\setminus\{i\}}x^{(t)}_{jm}(A_{i,j}-\beta)-2\alpha\sum_{j\in[n]\setminus\{i\}}\sum_{p\in[k]\setminus\{m\}}x^{(t)}_{jp}A_{i,j}}_{c_{im}}\big{)}+C.$ (40)
Therefore, we have
$G_{i,m}\triangleq[\nabla_{i}f({\bm{x}}^{(t)}_{[n]})]_{m}=c_{im},\quad\text{for}m\in[k].$ (41)
This holds because neither $c_{im}$ nor $C$ depend on $x^{(t)}_{im}$. Furthermore, since $x_{i0}$ does not show up in Eq. 40, everything in Eq. 40 is a constant w.r.t. $x_{i0}$. We therefore have
$G_{i,0}\triangleq[\nabla_{i}f({\bm{x}}^{(t)}_{[n]})]_{0}=0.$ (42)
By noting that ${\bm{x}}^{(t)}$ is discrete, we can rewrite $c_{im}$ for $m\in[k]$ as follows.
$c_{im}$ $=-\beta+2\sum_{j\in[n]\setminus\{i\}}x^{(t)}_{jm}(A_{i,j}-\beta)-2\alpha\sum_{j\in[n]\setminus\{i\}}\sum_{p\in[k]\setminus\{m\}}x^{(t)}_{jp}A_{i,j}$ (43)
$=-\beta+2\sum_{j\in[n]\setminus\{i\}}x^{(t)}_{jm}A_{i,j}-2\sum_{j\in[n]\setminus\{i\}}\beta-2\alpha\sum_{j\in[n]\setminus\{i\}}\sum_{p\in[k]\setminus\{m\}}x^{(t)}_{jp}A_{i,j}$
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2\sum_{j\in S_{m}\setminus\{i\}}\beta-2\alpha\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}$
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}.$
In the last equality we use $-2\sum_{j\in S_{m}\setminus\{i\}}\beta=-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}$. We note that computing the final expression in Eq. 43 is $O(k^{2}n)$, due to the last term. However, by noting that $\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}=\sum_{j\notin S_{0}}A_{i,j}-\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}$ we can derive the following.
$c_{im}$ $=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{p\in[k]\setminus\{m\}}\sum_{j\in S_{p}\setminus\{i\}}A_{i,j}$ (44)
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\big{(}\sum_{j\notin S_{0}}A_{i,j}-\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}\big{)}$
$=-\beta+2\sum_{j\in S_{m}\setminus\{i\}}(A_{i,j}+\alpha A_{i,j})-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{j\notin S_{0}}A_{i,j}$
$=-\beta+2(1+\alpha)\sum_{j\in S_{m}\setminus\{i\}}A_{i,j}-2|S_{m}|\beta+2\beta\bm{1}_{[i\in S_{m}]}-2\alpha\sum_{j\notin S_{0}}A_{i,j}$
Now, recall the defintions $M_{i,m}\triangleq 2\sum_{j\in S_{m}}A_{i,j}$, $\eta_{i}\triangleq\sum_{m\in[k]}M_{i,m}$ and $\beta_{im}=2\beta|S_{m}|-2\beta\bm{1}_{[i\in S_{m}]}$. Then, we note that $2\sum_{j\notin S_{0}}A_{i,j}=\sum_{m\in[k]}M_{i,m}$ (sum of all similarities from object $i$ to all non-neutral objects). Then, we obtain the final expression, i.e.,
$c_{im}=-\beta+(1+\alpha)M_{i,m}-\beta_{im}-\alpha\eta_{i},\quad\text{for }m\in[k],$ (45)
The expression in Eq. 45 can be computed in $O(kn)$ since it just amounts to computing $M_{i,m}$ for all $m\in[k]$ (i.e., the total similarity from $i$ to all non-neutral cluters $m\in[k]$).
See 4
Proof.
By examining the gradient in Eq. 45, we observe that the dominant term involving $\beta$ is $-\beta|S_{m}|$. Consequently, making $\beta$ large and negative increases the incentive to assign objects to non-neutral clusters. Moreover, since $-\beta|S_{m}|$ scales with cluster size, the local search procedure will favor placing an object $i$ in the largest non-neutral cluster. If $\beta$ is sufficiently large and negative, this term will completely dominate the objective, ensuring that no object is assigned to the neutral set (as the contribution to all non-neutral clusters remains positive). Ultimately, all objects will be placed in the largest non-neutral cluster.
Similarly, if $\beta$ is made very large and positive, $-\beta|S_{m}|$ will eventually dominate the objective, making the contribution to every non-neutral cluster negative for all objects. As a result, all objects will be assigned to the neutral set. ∎
Appendix B Experiments: More Details and Further Results
B.1 Datasets
Following [Tzeng et al., 2020], we consider the following widely studied real-world signed networks. WoW-EP8 (W8) [Kristof et al., 2020] represents interactions among authors in the 8th EU Parliament legislature, where edge signs indicate collaboration or competition. Bitcoin (BTC) [Leskovec and Krevl, 2014] is a trust-distrust network of users trading on the Bitcoin OTC platform. WikiVot (WikiV) [Leskovec and Krevl, 2014] records positive and negative votes for Wikipedia admin elections. Referendum (REF) [Lai et al., 2018] captures tweets about the 2016 Italian constitutional referendum, with edge signs indicating whether users share the same stance. Slashdot (SD) [Leskovec and Krevl, 2014] is a friend-foe network from the Slashdot Zoo feature. WikiCon (WikiC) [Kunegis, 2013] tracks positive and negative interactions between users editing English Wikipedia. Epinions (EP) [Leskovec and Krevl, 2014] represents the trust-distrust relationships in the Epinions online social network. WikiPol (WikiP) [Maniu et al., 2011] captures interactions among users editing Wikipedia pages on political topics.
B.2 Baselines
For SCG, we use the public implementation from [Ruo-Chun Tzeng, 2020]. For KOCG, we use the public implementation from [Lingyang Chu, 2016] with default hyperparameters: $\alpha=1/(k-1)$, $\beta=50$ (note that the purpose of this $\beta$ differs from the one used in our paper), and $\ell=5000$. For the spectral methods SPONGE and BNC, we use the public implementations from [Peter Davies and Aldo Glielmo, 2019]. Following [Tzeng et al., 2020], for SPONGE, we evaluate both the unnormalized and symmetric normalized versions and report results for the best-performing method.
B.3 Complete Results on All Datasets
Tables 4-5 report results for polarity and BA-Polarity with $k=2$, $k=4$, and $k=6$. The conclusions for $k=2$ are consistent with those for $k=4$ and $k=6$, discussed in the main paper. According to these results, our method consistently achieves the highest scores or performs competitively with the best scores.
Tables 6-13 present detailed analyses for all datasets across 12 different aspects, including polarity and BA-polarity. They provide insights into the differences between the clustering solutions computed by each method. In the context of PCD, evaluating clustering quality is inherently subjective, as different aspects capture different aspects of the solution. Generally, there is a trade-off among these aspects, and the objective is to achieve a good balance between them. Below, we define the aspects used in our analysis. We have defined these aspects such that a larger number is better (apart from runtime).
Let $N=\sum_{m\in[k]}|S_{m}|$ denote the number of non-neutral objects, and let $N_{\text{nz}}=N^{+}_{\text{intra}}+N^{-}_{\text{intra}}+N^{-}_{\text{inter}}+N^{+}_{\text{inter}}$ represent the number of non-zero similarities between non-neutral objects.
  • SIZE = $N$: The total number of non-neutral objects.
  • BAL = $S_{m}/S_{l}$: The balance factor, where $S_{m}$ and $S_{l}$ are the largest and smallest non-neutral clusters, respectively. Its range it $[0,1]$, with $1$ indicating perfect balance.
  • K: The number of non-empty non-neutral clusters.
  • MAC: Mean Average Cohesion, quantifying the density of positive intra-cluster similarities, defined as
    $\texttt{MAC}=\frac{1}{k}\sum_{m\in[k]}\frac{1}{|S_{m}|(|S_{m}|-1)}\sum_{i,j\in S_{m}}A^{+}_{i,j}.$
    Its range is $[0,1]$, where higher values indicate stronger cohesion within clusters.
  • MAO: Mean Average Opposition, measuring the density of negative inter-cluster similarities, defined as
    $\texttt{MAO}=\frac{1}{k(k-1)}\sum_{\begin{subarray}{c}m\in[k]\\ p\in[k]\setminus\{m\}\end{subarray}}\frac{1}{|S_{m}||S_{p}|}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A^{-}_{i,j}.$
    Its range is $[0,1]$, where higher values indicate stronger opposition between clusters.
  • CC+: Measures the fraction of intra-cluster similarities that are positive minus those that are negative, defined as
[…]
    Its range is $[-1,1]$, where $-1$ indicates that all non-zero intra-cluster similarities are negative, and $+1$ indicates that all are positive.
  • CC-: Measures the fraction of inter-cluster similarities that are negative minus those that are positive, defined as
[…]
    Its range is $[-1,1]$, where $-1$ indicates that all non-zero inter-cluster similarities are positive, and $+1$ indicates that all are negative.
  • DENS: The proportion of non-zero similarities among non-neutral objects, defined as
    $\texttt{DENS}=\frac{N_{\text{nz}}}{N(N-1)}.$
    Its range is $[0,1]$, with higher values indicating denser connectivity.
  • ISO: Isolation, measuring the separation between non-neutral and neutral objects, defined as
    $\texttt{ISO}=\frac{N_{\text{nz}}}{N_{\text{nz}}+\sum_{i\in S_{0}}\sum_{j\notin S_{0}}|A_{i,j}|}.$
    Its range is $[0,1]$, where ISO = $1$ means non-neutral objects are fully isolated from neutral ones, meaning no non-zero edges exist between them (which is ideal).
  • POL: Polarity, as defined in the main paper.
  • BA-POL: Balance-aware polarity, as defined in the main paper.
  • RT: Runtime of the corresponding method in seconds.
From Tables 6-13, we derive several key insights. Firstly, our method exhibits low standard deviation, indicating robustness to the initial random solution. It consistently finds high-polarity solutions while maintaining better balance than its main competitor, SCG. Unlike some baselines such as KOCG and SPONGE, our method does not enforce excessive balance, ensuring solutions remain both high in polarity and reasonably balanced.
In terms of runtime, our method is efficient and competitive with the baselines. It is also consistently ranked among the best in both DENS and ISO. Unlike SCG, our method always identifies $k$ non-empty non-neutral clusters, which we argue is a significant limitation of SCG.
While SCG generally achieves higher MAC values, this is largely due to its tendency to produce highly imbalanced solutions, often with singleton clusters. Since small or singleton clusters trivially yield high average cohesion (values close to 1), they can disproportionately inflate the overall MAC score.
Finally, our method performs comparably or better than SCG in CC+ and significantly outperforms it in CC- in most cases. This highlights another limitation of SCG: it often includes more positive similarities between clusters than negative ones, as reflected by negative CC- values. Some baselines produce either overly large or overly small non-neutral clusters (based on SIZE), whereas our method consistently finds solutions with a reasonable number of non-neutral objects (which consequently leads to a good balance of the other aspects), similar to SCG. However, we note that we can easily adjust the number of non-neutral objects by adjusting $\beta$, as discussed in the paper.
(a) Epinions, $k=2$, $\beta=0.1$
(b) Wikipol, $k=6$, $\alpha=\frac{1}{k-1}$
Figure 2: Investigation of the impact of $\alpha$ and $\beta$.
In Figure 2, we illustrate the impact of varying $\alpha$ and $\beta$. According to Figure 2(a), increasing $\alpha$ naturally balances intra-cluster cohesion and inter-cluster opposition. The size proportion, defined as the fraction of non-neutral objects in $V$, remains constant as $\alpha$ varies. Figure 2(b) shows that increasing $\beta$ monotonically reduces the number of non-neutral objects, leading to denser clusters, as indicated by improved MAC and MAO scores. Notably, balance remains stable across different $\beta$ values, unlike the baseline SCG-MA.
Table 4: Results for different methods and datasets w.r.t. the balance-aware polarity metric. $|E|$ is the number of non-zero edges and $|E_{-}|$ is the number of negative edges. We observe that our method (LSPCD) often yields the highest scores.
W8 BTC WikiV REF SD WikiC EP WikiP
$|V|$ 790 5.9K 7.1K 10.9K 82.1K 116.7K 131.6K 138.6K
$|E|$ 116K 214.5K 1M 251.4K 500.5K 2M 711.2K 715.9K
$|E_{-}|/|E|$ 0.2 0.2 0.2 0.1 0.2 0.6 0.2 0.1
$k=2$ LSPCD (ours) 19.4 10.9 13.9 61.8 9.7 102.2 56.2 12.4
SCG-MA 0.4 2.4 0.4 1.3 0.8 44.7 2.5 0.6
SCG-MO 0.5 0.6 0.4 1.3 0.7 39.6 2.4 0.4
SCG-B 4.9 19.7 0.7 1.7 1.5 47.3 3.8 0.9
SCG-R 3.4 1.8 4.6 2.1 1.1 31.9 6.6 3.0
KOCG-top-$1$ 10.4 1.0 3.8 4.8 1.3 3.5 3.3 1.8
KOCG-top-$r$ 11.3 3.3 2.2 11.7 2.2 2.9 10.1 1.1
BNC-$(k+1)$0.4 -0.9 -0.6 -1.0
BNC-$k$ 0.7 0.0 0.0 0.0
SPONGE-$(k+1)$ 19.7 0.6 0.3 0.6
SPONGE-$k$ 22.4 0.0 0.0 0.0
$k=4$ LSPCD (ours) 6.9 3.6 5.9 3.9 6.7 26.3 14.9 5.4
SCG-MA 5.8 0.3 0.8 0.3 1.9 0.7 0.1 0.7
SCG-MO 4.9 0.3 0.4 0.4 0.2 6.6 0.1 0.0
SCG-B 0.3 0.1 0.2 0.1 0.5 18.5 0.2 0.1
SCG-R 1.3 0.9 0.9 0.6 2.4 2.0 0.3 1.0
KOCG-top-$1$ 5.8 4.2 1.2 2.6 0.9 0.4 4.0 0.9
KOCG-top-$r$ 5.4 2.7 2.6 1.6 1.0 2.8 5.9 1.2
BNC-$(k+1)$0.1 -0.8 -0.4 -1.0
BNC-$k$ 0.5 0.0 0.0 0.0
SPONGE-$(k+1)$ 22.9 0.0 0.4 0.3
SPONGE-$k$ 18.6 0.0 0.0 0.0
$k=6$ LSPCD (ours) 4.4 2.2 3.6 2.2 3.8 16.4 6.8 3.5
SCG-MA 3.1 0.1 0.2 0.1 0.0 0.6 0.4 0.0
SCG-MO 3.6 0.1 0.2 0.1 0.1 4.2 0.1 0.0
SCG-B 0.3 0.2 0.2 0.1 0.2 1.8 0.2 0.1
SCG-R 0.3 0.3 0.4 0.2 0.4 2.3 0.1 0.0
KOCG-top-$1$ 6.1 1.7 2.5 3.3 1.7 0.4 2.7 3.7
KOCG-top-$r$ 5.0 1.2 1.8 2.5 1.3 1.1 2.8 0.7
BNC-$(k+1)$0.1 -0.1 -0.8 -0.4
BNC-$k$ 0.5 0.0 0.0 0.0
SPONGE-$(k+1)$ 17.7 0.0 0.4 0.6
SPONGE-$k$ 16.0 0.0 0.0 0.0
Table 5: Results for different methods and datasets w.r.t. the polarity metric. $|E|$ is the number of non-zero edges and $|E_{-}|$ is the number of negative edges. Our method (LSPCD) yields the best or competitive results, despite the fact that it, unlike SCG methods, does not directly optimize polarity.
W8 BTC WikiV REF SD WikiC EP WikiP
$|V|$ 790 5.9K 7.1K 10.9K 82.1K 116.7K 131.6K 138.6K
$|E|$ 116K 214.5K 1M 251.4K 500.5K 2M 711.2K 715.9K
$|E_{-}|/|E|$ 0.2 0.2 0.2 0.1 0.2 0.6 0.2 0.1
$k=2$ LSPCD (ours) 223.4 29.0 62.3 146.1 75.9 190.8 127.8 82.0
SCG-MA 236.6 28.8 71.5 172.2 77.5 155.2 128.3 82.8
SCG-MO 236.6 29.5 71.7 174.1 79.7 175.7 128.7 88.4
SCG-B 200.6 21.6 37.6 116.3 61.0 129.3 156.4 46.5
SCG-R 214.6 14.2 54.7 120.9 29.7 101.1 72.3 36.1
KOCG-top-$1$ 13.0 1.0 7.6 11.6 2.0 5.9 8.2 3.0
KOCG-top-$r$ 13.0 3.8 2.3 15.4 2.6 3.4 14.0 1.3
BNC-$(k+1)$0.7 -10.8 -1.1 -1.0
BNC-$k$ 184.6 5.3 15.8 41.5
SPONGE-$(k+1)$ 88.0 1.0 1.0 1.0
SPONGE-$k$ 191.4 5.1 15.8 41.5
$k=4$ LSPCD (ours) 218.5 23.3 52.6 139.2 61.1 113.6 111.5 71.6
SCG-MA 205.1 25.1 52.9 94.5 35.5 104.9 127.4 56.5
SCG-MO 213.2 25.3 53.1 82.1 38.5 117.9 129.0 39.7
SCG-B 211.6 12.4 24.8 116.2 48.3 49.8 94.4 45.7
SCG-R 214.6 8.0 19.5 118.7 10.7 41.1 65.1 33.7
KOCG-top-$1$ 9.1 8.4 4.5 15.0 2.6 4.5 8.9 3.1
KOCG-top-$r$ 7.4 5.0 3.3 3.7 3.0 3.8 11.0 4.4
BNC-$(k+1)$0.2 -9.4 -1.1 -1.0
BNC-$k$ 185.3 5.2 15.8 41.5
SPONGE-$(k+1)$ 53.8 1.1 1.0 1.0
SPONGE-$k$ 71.2 5.1 15.8 41.5
$k=6$ LSPCD (ours) 217.3 20.0 46.2 137.6 57.1 96.1 103.4 58.7
SCG-MA 207.3 14.6 45.5 84.9 37.8 102.6 88.8 57.5
SCG-MO 205.8 15.2 47.0 55.6 34.6 111.6 129.2 41.8
SCG-B 211.6 9.3 23.3 116.2 47.7 46.1 94.5 46.0
SCG-R 201.2 6.9 10.4 50.3 7.9 18.3 43.3 3.3
KOCG-top-$1$ 7.9 4.1 4.5 8.6 3.6 4.9 6.0 10.1
KOCG-top-$r$ 9.1 3.6 3.1 4.0 3.3 1.5 6.8 3.6
BNC-$(k+1)$0.2 -4.2 -1.1 -0.8
BNC-$k$ 185.2 5.2 15.8 41.5
SPONGE-$(k+1)$ 47.8 1.3 1.0 1.0
SPONGE-$k$ 57.9 5.1 15.8 41.5
Table 6: Detailed results on the WoW-EP8 dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 586 0.087 2 0.261 0.098 0.757 0.509 0.511 0.769 223.406 19.409 0.249
LSPCD (std) 0 0.0 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.034
SCG-MA 527 0.002 1 0.52 0.0 0.762 0.0 0.59 0.725 236.55 0.448 1.025
SCG-MO 517 0.002 1 0.527 0.0 0.769 0.0 0.596 0.708 236.592 0.457 1.026
SCG-B 583 0.025 2 0.274 0.094 0.697 -0.369 0.514 0.767 200.604 4.918 4.657
SCG-R 513 0.016 2 0.272 0.104 0.761 0.451 0.552 0.654 214.616 3.386 0.517
KOCG-top-$1$ 16 0.8 2 0.986 0.889 0.965 0.778 1.0 0.019 13.0 10.4
KOCG-top-$r$ 527 0.869 2 0.467 0.11 0.658 -0.598 0.559 0.691 12.964 11.269
BNC-$(k+1)$ 3 0.667 2 0.5 0.0 -1.0 0.0 0.333 0.007 -0.667 -0.444 0.99
BNC-$k$ 790 0.004 2 0.152 0.049 0.628 1.0 0.372 1.0 184.63 0.702 0.537
SPONGE-$(k+1)$ 375 0.224 2 0.204 0.095 0.738 0.318 0.343 0.346 87.957 19.705 0.581
SPONGE-$k$ 790 0.117 2 0.191 0.093 0.696 0.15 0.372 1.0 191.38 22.404 0.572
$k=4$ LSPCD (avg) 590 0.031 4 0.156 0.082 0.76 0.426 0.506 0.772 218.458 6.862 0.317
LSPCD (std) 1 0.0 0 0.003 0.002 0.001 0.016 0.0 0.002 0.142 0.008 0.044
SCG-MA 599 0.028 4 0.58 0.13 0.762 -0.32 0.527 0.822 205.137 5.828 1.198
SCG-MO 568 0.023 4 0.827 0.141 0.77 -0.349 0.55 0.776 213.211 4.939 1.176
SCG-B 615 0.002 1 0.421 0.0 0.693 0.0 0.498 0.822 211.561 0.343 6.233
SCG-R 503 0.006 4 0.211 0.056 0.762 0.747 0.564 0.644 214.623 1.301 1.131
KOCG-top-$1$ 31 0.636 4 0.962 0.668 0.944 0.339 0.978 0.039 9.054 5.761
KOCG-top-$r$ 599 0.735 4 0.436 0.099 0.692 -0.618 0.516 0.806 7.393 5.434
BNC-$(k+1)$ 8 0.4 4 0.5 0.0 -1.0 0.0 0.036 0.003 -0.25 -0.1 0.62
BNC-$k$ 790 0.003 4 0.327 0.03 0.632 1.0 0.372 1.0 185.305 0.473 0.594
SPONGE-$(k+1)$ 485 0.426 4 0.448 0.06 0.886 -0.53 0.33 0.433 53.823 22.911 0.593
SPONGE-$k$ 790 0.261 4 0.383 0.08 0.803 -0.499 0.372 1.0 71.162 18.585 0.61
$k=6$ LSPCD (avg) 591 0.02 6 0.142 0.071 0.761 0.414 0.505 0.773 217.344 4.419 0.269
LSPCD (std) 0 0.0 0 0.006 0.002 0.0 0.009 0.0 0.001 0.142 0.003 0.015
SCG-MA 598 0.015 6 0.785 0.113 0.763 -0.339 0.527 0.819 207.299 3.141 1.226
SCG-MO 591 0.017 6 0.756 0.171 0.77 -0.321 0.534 0.811 205.796 3.576 1.378
SCG-B 615 0.002 1 0.421 0.0 0.693 0.0 0.498 0.822 211.561 0.343 7.512
SCG-R 744 0.001 5 0.323 0.178 0.669 0.313 0.41 0.978 201.172 0.275 1.295
KOCG-top-$1$ 42 0.778 6 0.992 0.605 0.984 0.303 0.934 0.053 7.905 6.148
KOCG-top-$r$ 598 0.554 6 0.472 0.095 0.73 -0.637 0.522 0.812 9.109 5.045
BNC-$(k+1)$ 10 0.5 6 0.5 0.0 -1.0 0.0 0.022 0.002 -0.2 -0.1 0.977
BNC-$k$ 790 0.003 6 0.552 0.016 0.632 1.0 0.372 1.0 185.198 0.473 0.61
SPONGE-$(k+1)$ 572 0.372 6 0.537 0.065 0.893 -0.511 0.326 0.536 47.762 17.749 0.607
SPONGE-$k$ 790 0.276 6 0.53 0.075 0.834 -0.529 0.372 1.0 57.868 15.99 0.599
Table 7: Detailed results on the Bitcoin dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 155 0.377 2 0.2 0.143 0.94 0.969 0.199 0.211 29.022 10.947 2.203
LSPCD (std) 0 0.0 0 0.0 0.0 0.003 0.001 0.001 0.001 0.013 0.005 0.161
SCG-MA 179 0.084 2 0.298 0.068 0.906 0.873 0.179 0.216 28.838 2.418 0.122
SCG-MO 138 0.022 2 0.114 0.147 0.91 0.778 0.237 0.184 29.522 0.646 0.123
SCG-B 40 0.909 2 0.248 0.87 0.956 1.0 0.56 0.201 21.65 19.682 2.243
SCG-R 842 0.124 2 0.022 0.009 0.908 0.812 0.019 0.394 14.24 1.763 1.024
KOCG-top-$1$ 2 1.0 2 1.0 1.0 0.0 1.0 1.0 0.004 1.0 1.0
KOCG-top-$r$ 179 0.885 2 0.063 0.055 0.266 0.182 0.094 0.165 3.754 3.324
BNC-$(k+1)$ 50 0.083 2 0.058 0.0 -0.516 0.0 0.425 0.581 -10.76 -0.897 0.341
BNC-$k$ 5881 0.008 2 0.059 0.001 0.721 0.694 0.001 1.0 5.268 0.043 0.202
SPONGE-$(k+1)$ 6 0.6 2 0.667 0.0 1.0 0.0 0.2 1.0 1.0 0.6 1.116
SPONGE-$k$ 5881 0.001 2 0.501 0.0 0.697 0.0 0.001 1.0 5.092 0.003 2.138
$k=4$ LSPCD (avg) 217 0.154 4 0.182 0.12 0.929 0.815 0.143 0.256 23.333 3.586 2.117
LSPCD (std) 13 0.028 0 0.027 0.029 0.004 0.123 0.011 0.009 0.765 0.534 0.408
SCG-MA 176 0.014 4 0.488 0.07 0.914 -0.44 0.172 0.2 25.121 0.349 0.302
SCG-MO 180 0.014 4 0.487 0.07 0.91 -0.431 0.169 0.205 25.252 0.341 0.348
SCG-B 216 0.011 4 0.473 0.052 0.865 0.296 0.076 0.176 12.401 0.142 6.377
SCG-R 450 0.11 4 0.036 0.008 0.92 -0.627 0.033 0.238 8.033 0.886 1.237
KOCG-top-$1$ 26 0.5 4 0.859 0.621 1.0 0.653 0.738 0.112 8.41 4.205
KOCG-top-$r$ 176 0.54 4 0.136 0.041 0.856 -0.246 0.113 0.157 5.034 2.717
BNC-$(k+1)$ 58 0.083 4 0.112 0.0 -0.516 0.0 0.32 0.576 -9.414 -0.784 0.185
BNC-$k$ 5881 0.001 4 0.029 0.0 0.721 0.67 0.001 1.0 5.208 0.004 0.178
SPONGE-$(k+1)$ 71 0.045 4 0.754 0.0 1.0 0.0 0.016 0.443 1.099 0.05 3.364
SPONGE-$k$ 5881 0.001 4 0.75 0.0 0.697 0.0 0.001 1.0 5.092 0.003 2.797
$k=6$ LSPCD (avg) 194 0.114 6 0.251 0.143 0.948 0.646 0.155 0.231 20.031 2.201 2.73
LSPCD (std) 23 0.044 0 0.046 0.053 0.007 0.304 0.019 0.015 1.827 0.582 0.468
SCG-MA 430 0.009 6 0.536 0.021 0.931 -0.355 0.055 0.301 14.568 0.125 0.448
SCG-MO 412 0.009 6 0.571 0.028 0.929 -0.337 0.058 0.3 15.165 0.14 0.477
SCG-B 326 0.017 6 0.313 0.009 0.866 -0.421 0.053 0.222 9.321 0.16 10.509
SCG-R 860 0.038 6 0.038 0.006 0.941 -0.529 0.017 0.367 6.861 0.258 2.125
KOCG-top-$1$ 28 0.429 6 0.867 0.338 1.0 0.197 0.537 0.055 4.071 1.745
KOCG-top-$r$ 430 0.333 6 0.077 0.013 0.88 -0.405 0.043 0.26 3.601 1.2
BNC-$(k+1)$ 224 0.018 6 0.075 0.001 -0.622 0.958 0.033 0.394 -4.239 -0.077 0.286
BNC-$k$ 5881 0.001 6 0.075 0.0 0.722 0.657 0.001 1.0 5.197 0.003 0.194
SPONGE-$(k+1)$ 222 0.016 6 0.622 0.0 1.0 0.0 0.006 0.401 1.252 0.02 1.959
SPONGE-$k$ 5881 0.001 6 0.563 0.0 0.696 -1.0 0.001 1.0 5.085 0.003 2.473
Table 8: Detailed results on the WikiVot dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 1278 0.224 2 0.038 0.015 0.831 0.673 0.06 0.548 62.322 13.948 2.457
LSPCD (std) 4 0.001 0 0.0 0.0 0.0 0.003 0.0 0.001 0.003 0.056 0.276
SCG-MA 813 0.005 2 0.048 0.079 0.846 0.671 0.104 0.436 71.476 0.353 1.537
SCG-MO 748 0.005 2 0.052 0.082 0.854 0.671 0.113 0.411 71.733 0.385 1.432
SCG-B 414 0.02 2 0.054 0.033 0.756 0.776 0.12 0.221 37.589 0.737 7.501
SCG-R 1100 0.085 2 0.032 0.013 0.83 0.618 0.061 0.444 54.693 4.63 0.781
KOCG-top-$1$ 10 0.5 2 0.905 0.857 1.0 1.0 0.844 0.012 7.6 3.8
KOCG-top-$r$ 813 0.964 2 0.047 0.022 0.427 -0.337 0.066 0.297 2.312 2.229
BNC-$(k+1)$ 9 0.571 2 0.0 0.0 -1.0 0.0 0.139 1.0 -$1$0.635 0.721
BNC-$k$ 7115 0.002 2 0.002 0.0 0.558 0.0 0.004 1.0 15.794 0.024 0.49
SPONGE-$(k+1)$ 10 0.333 2 0.571 0.0 1.0 0.0 0.111 1.0 1.0 0.333 1.602
SPONGE-$k$ 7115 0.002 2 0.057 0.0 0.558 0.0 0.004 1.0 15.794 0.024 0.977
$k=4$ LSPCD (avg) 1089 0.115 4 0.073 0.013 0.856 -0.045 0.072 0.489 52.605 5.869 4.381
LSPCD (std) 149 0.029 0 0.026 0.002 0.004 0.427 0.011 0.04 6.003 0.86 1.69
SCG-MA 1142 0.015 4 0.081 0.018 0.849 -0.618 0.069 0.506 52.945 0.78 2.042
SCG-MO 1059 0.007 4 0.089 0.022 0.858 -0.692 0.073 0.474 53.07 0.356 1.986
SCG-B 790 0.01 4 0.091 0.014 0.774 -0.718 0.077 0.342 24.782 0.243 18.286
SCG-R 1524 0.044 4 0.031 0.008 0.813 -0.68 0.043 0.549 19.524 0.861 3.074
KOCG-top-$1$ 33 0.267 4 0.845 0.086 0.933 -0.609 0.576 0.03 4.525 1.207
KOCG-top-$r$ 1142 0.803 4 0.055 0.011 0.719 -0.618 0.059 0.44 3.288 2.639
BNC-$(k+1)$ 15 0.333 4 0.0 0.0 -1.0 0.0 0.076 1.0 -1.067 -0.356 0.527
BNC-$k$ 7115 0.0 4 0.001 0.0 0.558 0.0 0.004 1.0 15.794 0.007 0.533
SPONGE-$(k+1)$ 12 0.429 4 0.8 0.0 1.0 0.0 0.091 1.0 1.0 0.429 2.327
SPONGE-$k$ 7115 0.001 4 0.156 0.0 0.558 0.0 0.004 1.0 15.794 0.011 1.522
$k=6$ LSPCD (avg) 534 0.078 6 0.143 0.029 0.896 -0.314 0.133 0.287 46.179 3.586 5.292
LSPCD (std) 46 0.008 0 0.015 0.002 0.004 0.07 0.012 0.008 2.177 0.201 0.931
SCG-MA 1355 0.004 6 0.064 0.023 0.849 -0.647 0.056 0.564 45.494 0.168 2.24
SCG-MO 1226 0.004 6 0.073 0.024 0.859 -0.683 0.063 0.526 47.013 0.189 2.178
SCG-B 941 0.01 6 0.121 0.018 0.78 -0.735 0.065 0.369 23.332 0.229 29.072
SCG-R 1501 0.043 6 0.039 0.008 0.817 -0.734 0.044 0.542 10.433 0.444 3.475
KOCG-top-$1$ 40 0.556 6 0.894 0.227 0.981 -0.188 0.564 0.033 4.52 2.511
KOCG-top-$r$ 1355 0.561 6 0.051 0.009 0.73 -0.62 0.05 0.506 3.132 1.757
BNC-$(k+1)$ 13 0.75 6 0.0 0.0 -1.0 0.0 0.09 1.0 -1.077 -0.808 0.966
BNC-$k$ 7115 0.0 6 0.001 0.0 0.558 0.0 0.004 1.0 15.794 0.007 0.546
SPONGE-$(k+1)$ 20 0.429 6 0.644 0.0 1.0 0.0 0.053 1.0 1.0 0.429 1.791
SPONGE-$k$ 7115 0.0 6 0.434 0.0 0.558 0.0 0.004 1.0 15.794 0.007 1.66
Table 9: Detailed results on the Referendum dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 915 0.423 2 0.279 0.014 1.0 0.114 0.17 0.353 146.109 61.846 3.376
LSPCD (std) 1 0.002 0 0.001 0.0 0.0 0.017 0.001 0.001 0.092 0.233 0.19
SCG-MA 824 0.007 2 0.455 0.247 1.0 0.558 0.211 0.409 172.206 1.26 1.863
SCG-MO 673 0.007 2 0.546 0.3 1.0 0.571 0.261 0.352 174.083 1.299 1.108
SCG-B 1158 0.015 2 0.176 0.068 1.0 0.58 0.101 0.396 116.252 1.729 23.313
SCG-R 1550 0.018 2 0.095 0.04 1.0 0.529 0.079 0.492 120.85 2.14 4.657
KOCG-top-$1$ 15 0.417 2 0.973 0.659 1.0 1.0 0.829 0.007 11.6 4.833
KOCG-top-$r$ 824 0.761 2 0.057 0.018 0.705 -0.317 0.065 0.169 15.425 11.741
BNC-$(k+1)$ 4 1.0 2 0.0 0.0 -1.0 0.0 0.333 0.286 -1.0 -1.0 1.929
BNC-$k$ 10884 0.0 2 0.002 0.0 0.898 -1.0 0.004 1.0 41.495 0.011 1.114
SPONGE-$(k+1)$ 6 0.6 2 0.667 0.0 1.0 0.0 0.2 1.0 1.0 0.6 6.754
SPONGE-$k$ 10884 0.0 2 0.502 0.0 0.898 0.0 0.004 1.0 41.495 0.011 6.889
$k=4$ LSPCD (avg) 1065 0.028 4 0.196 0.043 1.0 0.056 0.145 0.394 139.163 3.915 3.724
LSPCD (std) 1 0.0 0 0.0 0.0 0.0 0.001 0.0 0.0 0.037 0.006 0.392
SCG-MA 1713 0.004 4 0.124 0.048 1.0 -0.693 0.081 0.512 94.544 0.348 6.809
SCG-MO 1658 0.004 4 0.142 0.054 1.0 -0.767 0.084 0.502 82.139 0.355 3.863
SCG-B 1142 0.001 1 0.102 0.0 1.0 0.0 0.102 0.398 116.233 0.102 60.02
SCG-R 1514 0.005 4 0.174 0.019 1.0 0.432 0.08 0.479 118.706 0.559 2.545
KOCG-top-$1$ 53 0.172 4 0.85 0.297 1.0 -0.363 0.615 0.024 14.956 2.579
KOCG-top-$r$ 1713 0.426 4 0.065 0.003 0.885 -0.862 0.052 0.363 3.711 1.581
BNC-$(k+1)$ 8 1.0 4 0.0 0.0 -1.0 0.0 0.143 0.25 -1.0 -1.0 1.125
BNC-$k$ 10884 0.0 4 0.001 0.0 0.898 -0.429 0.004 1.0 41.495 0.011 1.129
SPONGE-$(k+1)$ 18 0.333 4 0.452 0.0 1.0 0.0 0.059 1.0 1.0 0.333 5.042
SPONGE-$k$ 10884 0.0 4 0.156 0.0 0.898 0.0 0.004 1.0 41.495 0.019 6.327
$k=6$ LSPCD (avg) 1021 0.016 6 0.176 0.028 1.0 0.04 0.15 0.379 137.627 2.211 5.461
LSPCD (std) 1 0.001 0 0.001 0.0 0.0 0.001 0.0 0.0 0.131 0.081 2.813
SCG-MA 1945 0.001 5 0.107 0.033 1.0 -0.771 0.069 0.56 84.933 0.104 8.225
SCG-MO 2469 0.001 5 0.16 0.003 1.0 -0.853 0.049 0.629 55.571 0.071 6.925
SCG-B 1142 0.001 1 0.102 0.0 1.0 0.0 0.102 0.398 116.233 0.102 98.669
SCG-R 1660 0.005 6 0.08 0.038 0.986 -0.756 0.052 0.356 50.258 0.247 5.075
KOCG-top-$1$ 81 0.381 6 0.923 0.088 1.0 -0.673 0.536 0.032 8.622 3.285
KOCG-top-$r$ 1945 0.63 6 0.061 0.003 0.917 -0.876 0.053 0.442 4.037 2.543
BNC-$(k+1)$ 12 0.5 6 0.222 0.0 -0.714 0.0 0.106 0.25 -0.833 -0.417 1.923
BNC-$k$ 10884 0.0 6 0.056 0.0 0.898 -0.2 0.004 1.0 41.495 0.011 1.155
SPONGE-$(k+1)$ 18 0.6 6 0.667 0.0 1.0 0.0 0.059 1.0 1.0 0.6 11.664
SPONGE-$k$ 10884 0.0 6 0.501 0.0 0.898 0.0 0.004 1.0 41.495 0.011 9.346
Table 10: Detailed results on the Slashdot dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 235 0.128 2 0.207 0.055 0.969 0.836 0.337 0.167 75.903 9.686 29.957
LSPCD (std) 0 0.002 0 0.0 0.001 0.0 0.003 0.001 0.0 0.095 0.133 3.151
SCG-MA 307 0.01 2 0.63 0.123 0.968 0.923 0.262 0.152 77.485 0.76 3.316
SCG-MO 234 0.009 2 0.674 0.137 0.973 0.882 0.352 0.145 79.692 0.681 2.654
SCG-B 289 0.025 2 0.145 0.056 0.98 -0.005 0.221 0.205 60.962 1.503 287.233
SCG-R 3033 0.036 2 0.007 0.007 0.872 0.635 0.011 0.216 29.706 1.065 25.778
KOCG-top-$1$ 3 0.667 2 1.0 1.0 1.0 1.0 1.0 0.005 2.0 1.333
KOCG-top-$r$ 307 0.828 2 0.028 0.03 0.159 0.182 0.05 0.037 2.612 2.164
$k=4$ LSPCD (avg) 380 0.11 4 0.212 0.087 0.966 0.492 0.192 0.189 61.089 6.738 37.751
LSPCD (std) 2 0.005 0 0.01 0.003 0.0 0.004 0.002 0.001 0.251 0.261 3.013
SCG-MA 2552 0.054 4 0.159 0.012 0.862 -0.431 0.026 0.269 35.53 1.907 16.032
SCG-MO 2111 0.004 4 0.181 0.051 0.876 -0.657 0.03 0.24 38.534 0.153 21.868
SCG-B 410 0.011 4 0.287 0.101 0.973 -0.491 0.128 0.199 48.306 0.51 814.654
SCG-R 3853 0.223 4 0.01 0.002 0.877 -0.34 0.008 0.227 10.749 2.396 27.234
KOCG-top-$1$ 23 0.364 4 0.453 0.172 1.0 0.643 0.206 0.009 2.609 0.949
KOCG-top-$r$ 2552 0.34 4 0.013 0.003 0.627 -0.477 0.012 0.16 2.973 1.012
$k=6$ LSPCD (avg) 272 0.066 6 0.306 0.083 0.982 0.423 0.251 0.156 57.075 3.759 64.95
LSPCD (std) 30 0.009 0 0.041 0.013 0.001 0.134 0.032 0.014 2.428 0.393 24.494
SCG-MA 2343 0.001 5 0.35 0.026 0.868 -0.701 0.028 0.256 37.849 0.02 32.081
SCG-MO 2504 0.003 6 0.212 0.063 0.876 -0.421 0.026 0.265 34.649 0.098 26.278
SCG-B 420 0.004 3 0.254 0.005 0.971 -0.481 0.124 0.191 47.676 0.168 1408.849
SCG-R 9661 0.045 6 0.02 0.002 0.814 -0.43 0.004 0.433 7.906 0.359 73.943
KOCG-top-$1$ 48 0.467 6 0.65 0.079 0.978 -0.166 0.216 0.016 3.583 1.672
KOCG-top-$r$ 2343 0.408 6 0.021 0.003 0.722 -0.54 0.014 0.164 3.28 1.34
Table 11: Detailed results on the WikiCon dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 1876 0.536 2 0.055 0.128 0.871 0.997 0.108 0.242 190.8 102.249 72.368
LSPCD (std) 0 0.0 0 0.0 0.0 0.0 0.0 0.0 0.0 0.019 0.066 17.051
SCG-MA 8903 0.288 2 0.008 0.026 0.81 0.998 0.019 0.473 155.215 44.697 69.196
SCG-MO 2442 0.226 2 0.036 0.094 0.839 0.999 0.08 0.22 175.654 39.641 19.316
SCG-B 502 0.366 2 0.117 0.387 0.816 0.926 0.295 0.142 129.335 47.318 1314.819
SCG-R 12669 0.316 2 0.004 0.011 0.798 0.997 0.009 0.441 101.138 31.924 169.997
KOCG-top-$1$ 14 0.6 2 0.803 0.289 0.85 0.368 0.648 0.002 5.857 3.514
KOCG-top-$r$ 8903 0.849 2 0.007 0.007 0.0 0.056 0.014 0.327 3.417 2.901
$k=4$ LSPCD (avg) 2288 0.232 4 0.033 0.051 0.869 0.936 0.086 0.23 113.637 26.258 88.288
LSPCD (std) 40 0.02 0 0.008 0.005 0.002 0.058 0.001 0.009 2.613 1.623 20.523
SCG-MA 4852 0.007 4 0.042 0.104 0.82 0.577 0.027 0.241 104.937 0.707 139.274
SCG-MO 1943 0.056 4 0.063 0.117 0.848 0.533 0.086 0.163 117.935 6.581 69.637
SCG-B 1700 0.372 4 0.12 0.032 0.768 0.268 0.07 0.174 49.824 18.514 3792.395
SCG-R 7174 0.048 4 0.006 0.012 0.836 0.308 0.015 0.293 41.125 1.984 131.226
KOCG-top-$1$ 57 0.085 4 0.708 0.502 0.75 0.891 0.253 0.027 4.456 0.379
KOCG-top-$r$ 4852 0.742 4 0.014 0.011 0.213 -0.075 0.024 0.199 3.821 2.833
$k=6$ LSPCD (avg) 2394 0.172 6 0.049 0.039 0.873 0.847 0.08 0.233 96.085 16.382 69.593
LSPCD (std) 207 0.031 0 0.025 0.007 0.004 0.112 0.006 0.019 4.787 2.244 9.767
SCG-MA 4827 0.006 6 0.009 0.044 0.821 0.622 0.028 0.243 102.611 0.578 145.295
SCG-MO 2016 0.037 6 0.06 0.06 0.848 0.685 0.079 0.159 111.578 4.151 76.205
SCG-B 1924 0.039 6 0.084 0.015 0.771 0.291 0.061 0.174 46.069 1.779 6125.694
SCG-R 12909 0.128 6 0.011 0.004 0.788 0.135 0.009 0.463 18.278 2.331 175.294
KOCG-top-$1$ 50 0.074 6 0.765 0.476 0.962 0.633 0.505 0.007 4.904 0.363
KOCG-top-$r$ 4827 0.698 6 0.016 0.009 0.286 -0.209 0.023 0.2 1.522 1.062
Table 12: Detailed results on the Epinions dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 2188 0.44 2 0.12 0.013 0.907 0.74 0.066 0.351 127.784 56.221 59.119
LSPCD (std) 4 0.004 0 0.001 0.0 0.002 0.002 0.0 0.0 0.181 0.42 3.692
SCG-MA 1234 0.02 2 0.088 0.114 0.906 0.739 0.116 0.246 128.316 2.541 34.752
SCG-MO 1017 0.019 2 0.099 0.138 0.91 0.713 0.14 0.22 128.722 2.446 25.471
SCG-B 253 0.024 2 0.419 0.205 0.999 1.0 0.621 0.501 156.379 3.768 822.236
SCG-R 4396 0.091 2 0.01 0.007 0.891 0.766 0.019 0.363 72.282 6.561 12.119
KOCG-top-$1$ 12 0.4 2 0.708 0.815 1.0 0.833 0.803 0.007 8.167 3.267
KOCG-top-$r$ 1234 0.719 2 0.054 0.022 0.5 -0.245 0.064 0.16 14.036 10.092
$k=4$ LSPCD (avg) 2120 0.136 4 0.124 0.016 0.932 0.408 0.065 0.341 111.544 14.851 65.489
LSPCD (std) 129 0.042 0 0.021 0.003 0.002 0.312 0.004 0.006 7.5 3.548 2.958
SCG-MA 1576 0.001 3 0.416 0.001 0.928 -0.714 0.09 0.285 127.432 0.107 42.784
SCG-MO 1373 0.001 3 0.438 0.001 0.934 -0.635 0.103 0.264 128.951 0.129 34.407
SCG-B 868 0.002 3 0.405 0.0 0.926 -0.411 0.119 0.226 94.43 0.187 2169.558
SCG-R 1872 0.005 4 0.152 0.033 0.928 -0.801 0.044 0.23 65.124 0.294 49.068
KOCG-top-$1$ 28 0.455 4 0.865 0.62 0.953 0.582 0.81 0.011 8.905 4.048
KOCG-top-$r$ 1576 0.533 4 0.071 0.01 0.768 -0.63 0.06 0.202 11.001 5.869
$k=6$ LSPCD (avg) 2660 0.066 6 0.088 0.014 0.929 0.324 0.05 0.373 103.375 6.835 107.579
LSPCD (std) 153 0.009 0 0.009 0.004 0.002 0.142 0.004 0.007 3.637 0.857 20.806
SCG-MA 2564 0.005 6 0.301 0.048 0.935 -0.713 0.05 0.34 88.759 0.448 57.185
SCG-MO 1373 0.001 3 0.438 0.001 0.934 -0.635 0.103 0.264 129.22 0.129 37.194
SCG-B 868 0.002 3 0.405 0.0 0.926 -0.411 0.119 0.226 94.476 0.187 3696.279
SCG-R 1365 0.003 6 0.128 0.036 0.946 -0.898 0.054 0.203 43.324 0.138 53.993
KOCG-top-$1$ 34 0.444 6 0.9 0.43 1.0 0.496 0.576 0.012 5.965 2.651
KOCG-top-$r$ 2564 0.405 6 0.043 0.006 0.779 -0.654 0.035 0.262 6.802 2.753
Table 13: Detailed results on the WikiPol dataset from various aspects. LSPCD (avg) and LSPCD (std) respectively indicate the mean and standard deviation across five runs of our method with different seeds.
SIZE BAL K MAC MAO CC+ CC- DENS ISO POL BA-POL RT
$k=2$ LSPCD (avg) 599 0.151 2 0.093 0.034 0.917 0.87 0.15 0.109 81.985 12.417 57.022
LSPCD (std) 2 0.001 0 0.001 0.0 0.003 0.011 0.0 0.0 0.037 0.042 7.851
SCG-MA 1251 0.007 2 0.035 0.054 0.924 0.928 0.072 0.172 82.822 0.599 11.484
SCG-MO 648 0.005 2 0.071 0.079 0.928 1.0 0.147 0.121 88.441 0.41 4.041
SCG-B 609 0.02 2 0.041 0.013 0.963 -0.238 0.081 0.112 46.525 0.932 773.37
SCG-R 7400 0.082 2 0.003 0.001 0.91 0.63 0.005 0.305 36.119 2.968 76.435
KOCG-top-$1$ 6 0.6 2 0.75 0.625 1.0 1.0 0.6 0.003 3.0 1.8
KOCG-top-$r$ 1251 0.859 2 0.024 0.012 0.322 -0.284 0.035 0.097 1.258 1.081
$k=4$ LSPCD (avg) 450 0.075 4 0.147 0.068 0.938 0.546 0.184 0.086 71.628 5.389 83.214
LSPCD (std) 23 0.002 0 0.063 0.032 0.002 0.249 0.014 0.003 2.015 0.189 26.749
SCG-MA 2140 0.013 4 0.093 0.014 0.917 -0.613 0.038 0.217 56.471 0.728 49.769
SCG-MO 2783 0.0 3 0.283 0.001 0.895 -0.775 0.026 0.242 39.698 0.019 44.39
SCG-B 727 0.002 2 0.203 0.001 0.967 -0.899 0.07 0.125 45.661 0.076 2225.353
SCG-R 7740 0.029 4 0.002 0.001 0.916 0.599 0.005 0.302 33.723 0.966 91.037
KOCG-top-$1$ 26 0.286 4 0.558 0.119 0.949 0.182 0.255 0.005 3.051 0.872
KOCG-top-$r$ 2140 0.27 4 0.02 0.002 0.808 -0.609 0.01 0.092 4.409 1.192
$k=6$ LSPCD (avg) 825 0.06 6 0.191 0.026 0.94 -0.28 0.093 0.123 58.694 3.532 228.675
LSPCD (std) 72 0.001 0 0.023 0.014 0.005 0.183 0.011 0.008 1.548 0.079 192.894
SCG-MA 2176 0.001 4 0.259 0.001 0.919 -0.73 0.037 0.22 57.546 0.046 54.104
SCG-MO 2783 0.0 3 0.283 0.001 0.895 -0.775 0.026 0.242 41.846 0.02 48.571
SCG-B 727 0.002 2 0.203 0.001 0.967 -0.899 0.07 0.125 45.986 0.077 3973.384
SCG-R 95033 0.002 6 0.006 0.0 0.884 -0.686 0.003 0.901 3.329 0.007 331.066
KOCG-top-$1$ 83 0.364 6 0.756 0.029 0.967 -0.658 0.268 0.015 10.135 3.685
KOCG-top-$r$ 2176 0.192 6 0.032 0.002 0.867 -0.578 0.006 0.077 3.585 0.687
Appendix C Limitations of Polarity
[…]
The polarity values for these solutions are: $\text{Polarity}(S^{(1)})=(20+10)/8=3.75$, $\text{Polarity}(S^{(2)})=(38+6)/12=3.67$, and $\text{Polarity}(S^{(3)})=(30+0)/8=3.75$.
Although $S^{(1)}$ and $S^{(3)}$ achieve the same polarity score, $S^{(1)}$ is significantly more balanced, making it the more reasonable choice. In contrast, evaluating our objective (Eq. 7) for the same solutions, we obtain $S^{(1)}:(20+10)-(4^{2}+4^{2})=-2$, $S^{(2)}:(38+6)-(4^{2}+8^{2})=-36$, and $S^{(3)}:(30+0)-8^{2}=-34$.
Our objective function still identifies $S^{(2)}$ as the worst solution (consistent with polarity), but it strongly favors $S^{(1)}$ over $S^{(3)}$ due to its better balance. Here, we assume $\alpha=1/(k-1)$ (which is 1 since $k=2$) for consistency with polarity, and $\beta=1$.